### v1

# Smeared phase transition in the dissipative random quantum Ashkin-Teller model

Abstract
We study the effects of dissipation in the phase diagram of the random quantum Ashkin-Teller model by means of a generalization of the strong-disorder renormalization group combined with adiabatic renormalization. This model has three phases and three quantum phase transitions. We demonstrate that the combined effect of Ohmic dissipation and quenched disorder smears two out of the three quantum phase transitions. Our analytical theory allows us to understand why one of the phase transitions remains sharp. This is due to a cancellation of the dissipation effects on the nontrivial nature of the intertwined order parameter of one of the phases.
I Introduction
Quenched disorder underlies a manifold of profound effects near quantum phase transitions: For instance, disorder-induced inhomogeneities lead to exotic universality classes such as those typified by the infinite randomness critical point (fisher_1; fisher_2), and the strong off-critical Griffiths singularities (griffiths_original; mccoy-prl69). Furthermore, in certain situations these inhomogeneities could lead to the destruction or smearing (smeared_hoyos_1; smeared_hoyos_2) of the putative quantum phase transition. The moorings of the effects described above is centered around the so-called rare regions (RRs): arbitrarily large regions in space which, due to a statistically rare configuration of the quenched-disordered couplings, is locally ordered and weakly coupled to the bulk. These RRs, characterized by slow dynamics, dominate the low-energy behavior of physical observables (glass_1; glass_2), (for reviews, see Refs. (vojta2006rare; igloi-review; vojta2013phases; igloi-monthus-review2)).
In practice it is important to realize that real materials are seldom described as closed systems. Thus, it is of paramount importance to understand how the phases and phase transitions gets affected when the system is connected to a dissipative reservoir. The combined effects of quenched disorder and dissipation have been investigated in the Griffiths phase where it was established that sufficiently strong dissipation can completely damp the fluctuations of a RR (millis-morr-schmalian-prl01; vojta-prl03; vojta-schmalian-prb05; vojta-hoyos-prl14). Further, Strong-Disorder Renormalization-Group (SDRG) based analytical study established that for the case of the Random Transverse-Field Ising Model (RTFIM), in any dimension, the critical point is entirely destroyed by smearing when dissipation is Ohmic or Subohmic (smeared_hoyos_1; smeared_hoyos_2). Physically, the smearing phenomena occurs because dissipation freezes the dynamics of sufficiently large RRs, independent of the bulk, i.e., each rare region orders independently from the bulk, leading to local phase transitions that can take place at different values of the control parameter. Thus, the paramagnetic Griffiths phase becomes an inhomogeneously ordered ferromagnetic phase. Notably, experimental evidence of smeared phase transitions has been observed in the compounds Sr1-xCaxRuO3 (experimental_1) and CePd1-xRhx (experimental_2). More recently, smeared phase transitions have also been identified in a numerical study of strange metals (strangemetals), and have also potentially seen in low-dimensional superconductors (kaur-etal-njp24).
This work belongs to the genre of problems that we have discussed in the previous paragraph: Namely, we study the dissipative random quantum Ashkin-Teller (RQAT) model. Besides the usual paramagnetic (PM) and ferromagnetic (FM) phase, this model features an additional phase, called the product phase, which exhibits a composite (or intertwined) order (1981_ashkinteller). We inquire whether the transitions between this product phase and the other phases are smeared by the combined effects of dissipation and disorder. Interestingly, we show that Ohmic dissipation smears only the transition between the product and FM phases, while the transition between product and PM phases remains sharp. This conclusion is robust and persists across different forms of system-bath couplings. Surprisingly, even though dissipation is coupled to the original order parameter fields, it does not couple generically to the intertwined order parameter field and, therefore, does not damp the quantum fluctuations of the associated rare regions that form near the transition line between the product and PM phases.
Finally, besides this fundamental motivation, it is worth mentioning that the Ashkin-Teller model has applications in other contexts such as to describe layers of atoms absorbed on surfaces (bak-etal-prl85), current loops in high-Tc superconductors (aji-varma-prl07), and the elastic response of DNA molecules (chang-wang-zheng-ctp08). The understanding of dissipation in these contexts is another motivation for our work.
In the remainder of this work, we review the effects of disorder on the quantum Ashkin-Teller chain in Sec. II, introduce the model studied in Sec. III, apply the adiabatic renormalization-group and the strong-disorder renormalization-group methods to our problem in Secs. IV and V, respectively. The resulting phase diagram is then discussed in Sec. VI. Finally, we give concluding remarks in Sec. VII. Other possible models of dissipation are considered in Appendix A and we find that our conclusions remain qualitatively correct.
II Review of the disspationless model
The Hamiltonian of the one-dimensional random quantum Ashkin-Teller (RQAT) model is (1943_ashkinteller; 1981_ashkinteller; 2001_ashkinteller; 2008_ashkinteller; 2014_ashkinteller)
$H_{\text{S}}=$ $-\sum_{\nu=1}^{2}\sum_{i}\left(J_{i}S_{\nu,i}^{z}S_{\nu,i+1}^{z}+h_{i}S_{\nu,i}^{x}\right)$
$-\sum_{i}\left(K_{i}S_{1,i}^{z}S_{1,i+1}^{z}S_{2,i}^{z}S_{2,i+1}^{z}+g_{i}S_{1,i}^{x}S_{2,i}^{x}\right),$ (1)
where $S_{\nu,i}^{z}$ and $S_{\nu,i}^{x}$ are Pauli matrices representing local spin-1/2 moments. The first two terms on the right-hand-side of Eq. (1) describe two independent random transverse-field Ising (RTFI) chains whereas the remaining terms quantify the couplings between them. The index $\nu=1,2$, referred to as the color index, distinguishes the two chains and the index $i$ distinguishes different lattice sites. The interactions $J_{i}$ and transverse fields $h_{i}$ are strictly positive, independent random variables. The coupling strengths are characterized by the ratios $\epsilon_{h,i}=g_{i}/h_{i}$ and $\epsilon_{J,i}=K_{i}/J_{i}$. For simplicity, we assume that the (bare) values $\epsilon_{h,i}=\epsilon_{J,i}=\epsilon_{I}>0$, i.e., they are uniform and site independent. However, they flow under the renormalization-group transformations and become random variables (2014_ashkinteller). The low-energy physical behavior of the model (1) can be divided into two regimes: the weak-coupling regime ($\epsilon_{I}<1$) and the strong-coupling regime ($\epsilon_{I}>1$). In the low-energy limit of the weak-coupling regime, the two Ising chains effectively decouple (2001_ashkinteller; 2008_ashkinteller) and the resulting physics is that of two identical RTFI chains. Therefore, there is a direct quantum phase transition between a ferromagnet (where the quantum average $\left\langle S_{\nu,i}^{z}\right\rangle\neq 0$) and paramagnet (where $\left\langle S_{\nu,i}^{z}\right\rangle=0$). The transition, which occurs when $\prod J_{i}=\prod h_{i}$, is of infinite-randomness type and is surrounded by the Griffiths ferro- and paramagnetic phases. In the strong-coupling regime $\epsilon_{I}>1$, on the other hand, this transition is through an intervening product phase exhibiting a composite order, i.e., $\left\langle S_{\nu,i}^{z}\right\rangle=0$, but $\left\langle S_{1,i}^{z}S_{2,i}^{z}\right\rangle\neq 0$. The transition between the ferromagnet and the product phase is also of infinite-randomness type surrounded by quantum Griffiths ferromagnetic and product phases. By duality, [^1] The Hamiltonian (1) is invariant under the duality transformation: $S_{\nu,i}^{z}S_{\nu,i+1}^{z}\rightarrow\tau_{\nu,i}^{x}$, $S_{\nu,i}^{x}\rightarrow\tau_{\nu,i}^{z}\tau_{\nu,i+1}^{z}$ (here, $\tau^{x}$ and $\tau^{z}$ are the dual Pauli operators), $J_{i}\leftrightarrow h_{i}$, and $K_{i}\leftrightarrow g_{i}$. the transition between the paramagnetic and product phases is in the same universality class (2014_ashkinteller).
This result is better illustrated by reformulating the problem in terms of “product” variables $\boldsymbol{\sigma}_{i}$ (which are also spin-1/2 operators represented by Pauli matrices) and one of the “original” spin variable $\boldsymbol{\eta}_{i}$, defined as
$\sigma_{i}^{z}=S_{1,i}^{z}S_{2,i}^{z},\;\eta_{i}^{z}=S_{1,i}^{z},\;\sigma_{i}^{z}\eta_{i}^{z}=S_{2,i}^{z},$ (2)
and
$\eta_{i}^{x}=S_{1,i}^{x}S_{2,i}^{x},\;\sigma_{i}^{x}=S_{2,i}^{x},\;\sigma_{i}^{x}\eta_{i}^{x}=S_{1,i}^{x}.$ (3)
In these new variables, the Hamiltonian (1) becomes
$H_{\text{S}}=$ $-\sum_{i}\left(K_{i}\sigma_{i}^{z}\sigma_{i+1}^{z}+h_{i}\sigma_{i}^{x}\right)-\sum_{i}\left(J_{i}\eta_{i}^{z}\eta_{i+1}^{z}+g_{i}\eta_{i}^{x}\right)$
$-\sum_{i}\left(J_{i}\sigma_{i}^{z}\sigma_{i+1}^{z}\eta_{i}^{z}\eta_{i+1}^{z}+h_{i}\sigma_{i}^{x}\eta_{i}^{x}\right).$ (4)
From the first and second sums of the Hamiltonian (4), it is clear that, when $\epsilon_{I}\gg 1$ and sufficiently near to the self-duality line $\left(\delta\approx 0\right)$, the system is in a phase with long-range order in the $\boldsymbol{\sigma}$ variables ($\left\langle\sigma^{z}\right\rangle\neq 0$) and disordered in the $\boldsymbol{\eta}$ variables ($\left\langle\eta^{z}\right\rangle=0$). This implies that long-range order develops in the product variables $\left\langle S_{1}^{z}S_{2}^{z}\right\rangle\neq 0$, while the spins remain disordered in each individual color $\left\langle S_{1}^{z}\right\rangle=\left\langle S_{2}^{z}\right\rangle=0$ (1943_ashkinteller; 1981_ashkinteller). Thus, the moniker product (PROD) phase. A caveat is needed here. The terminology $\boldsymbol{\sigma}_{i}$ and $\boldsymbol{\eta}_{i}$ as a product and original variables is due to the lack of a better one. It is precise if one considers only the $z$-component of the original spins [see Eq. (2)]. Regarding the $x$-component [see Eq. (3)], however, the terminology is the other way round. Regarding the $y$-component, both could be regarded as product variables. In sum, both $\boldsymbol{\sigma}_{i}$ and $\boldsymbol{\eta}_{i}$ are nontrivial compositions of the original spin variables. We stick with this terminology because the $z$-component is the one defining the order parameter field of the transition.
In the presence of quenched disorder, it was found that all quantum phase transitions belong to the same universality class (that of the random transverse-field Ising chain) and are surrounded by the associated quantum Griffiths phases (2014_ashkinteller) as illustrated in the phase diagram Fig. 1(a).
Figure 1: (a) Schematic phase diagram of the RQAT chain without dissipation. The dashed lines indicate crossovers from conventional FM, PM and product phases to Griffiths phases. GFM and GPM denote the Griffiths ferromagnetic and paramagnetic phases, respectively. DG corresponds to the double Griffiths phase, characterized by the presence of rare regions in both $\sigma$ and $\eta$ variables. GPROD denotes the Griffiths product phase, where rare regions form in the $\sigma$ variable near the PM transition and in the $\eta$ variable near the FM transition. (b) Schematic phase diagram of the RQAT chain with dissipation. The transitions where rare regions are formed in the original spin variable are smeared. In the strong-coupling regime, the GFM phase is replaced by a weakly ordered (WO) phase and the GPROD phase is replaced by an inhomogeneous ferromagnetic (IFM) phase.
It is important to distinguish between the two sorts of Griffiths phases in the region $\epsilon_{I}>1$. Surrounding the PM–PROD transition, the Griffiths singularities in both the Griffiths PM and Griffiths PROD phases are due to rare regions locally in a PROD phase and weakly interacting with the bulk. In RRs of this kind, the $z$-component of the original spins fluctuate rapidly and independently of the nearby spins. However, they fluctuate together with the other color spin in the same site. On the other hand, surrounding the PROD-FM transition, the low-energy behavior of both the Griffiths PROD and FM phases are dominated by RRs locally in the FM phase and weakly coupled to the bulk. Near the multicritical point, both RRs coexist in the region $\epsilon_{I}>1$ and, therefore, give rise to the double Griffiths (DG) phase (2014_ashkinteller). It is worth noting that this DG phase have been seen in a density-matrix renormalization-group study (chatelain-voliotis-epjb16).
Evidently, these two phase transitions and the associated Griffiths phases are related by duality. However, this will be broken when dissipation is introduced as it couples to the original spin variables and not to the product ones. As a result, only the FM-like RRs will be dramatically affected by dissipation yielding to the smearing of the transition as sketched in Fig. c.
III A simple dissipative model
We introduce dissipation by coupling each spin to its own set of identical bosonic baths. The Hamiltonian thus becomes
$H=H_{\text{S}}+\sum_{\nu=1}^{2}H_{\text{B}}^{(\nu)}+\sum_{\nu=1}^{2}H_{\text{SB}}^{(\nu)}$ (5)
where $H_{\text{S}}$ is the RQAT Hamiltonian defined in Eq. (1). The bath component are independent harmonic oscillators, and thus,
$H_{\text{B}}^{(\nu)}=\sum_{k,i}\left(\frac{p_{\nu;k,i}^{2}}{2m_{k,i}}+\frac{m_{k,i}\omega_{k,i}^{2}x_{\nu;k,i}^{2}}{2}\right).$ (6)
Here, the oscillator frequencies $\omega_{k,i}$ and masses $m_{k,i}$ are assumed to be color-independent. Finally, following the Caldeira-Leggett model (caldeiraleggett; leggett_spinboson; smeared_hoyos_1; smeared_hoyos_2), the system-bath coupling is
$H_{\text{SB}}^{(\nu)}=\sum_{i}S_{\nu,i}^{z}\sum_{k}\lambda_{k,i}x_{\nu;k,i},$ (7)
where the constant $\lambda_{k,i}$ controls the strength of the coupling. For the moment, we consider color-independent coupling constants. Later, we discuss the effects of color-dependence.
The relevant properties of the baths are described by their spectral densities $F_{i}(\omega)$, defined as
$F_{i}(\omega)=\frac{\pi}{2}\sum_{k}\frac{\lambda_{k,i}^{2}}{m_{k,i}\omega_{k,i}}\delta(\omega-\omega_{k,i})=\frac{\pi}{2}\alpha_{i}\frac{\omega^{s}}{\omega_{c}^{s-1}}e^{-\omega/\omega_{c}}.$ (8)
Here, $\omega_{c}$ is the cutoff frequency, assumed to be identical for all sites, and $\alpha_{i}$ is a dimensionless parameter quantifying the dissipation strength. The exponent $s$ defines the dissipation regime: $s=1$, >1>1, and <1<1 corresponding to the Ohmic, super-Ohmic, and sub-Ohmic regimes, respectively.
Since we are interested in the $\epsilon>1$ regime, it is convenient to reformulate the Hamiltonian using the product variables defined in Eqs. (2) and (3). The system-bath Hamiltonian becomes
$H_{\text{SB}}^{(1)}+H_{\text{SB}}^{(2)}=\sum_{i}\eta_{i}^{z}\sum_{k}\lambda_{k,i}x_{1;k,i}+\sum_{i}\sigma_{i}^{z}\eta_{i}^{z}\sum_{k}\lambda_{k,i}x_{2;k,i}.$ (9)
Notice that the original variable $\boldsymbol{\eta}$ and the product variable $\boldsymbol{\sigma}$ do not couple equally to the oscillator baths. We will show that this has a profound consequence: the effects of dissipation on the PM-PROD and PROD-FM phase transitions are fundamentally different.
In what follows, we follow Refs. (smeared_hoyos_1; smeared_hoyos_2) to generalize the adiabatic renormalization procedure (leggett_spinboson) to the present case. Importantly, we assume weak dissipation ($\alpha_{i}\ll 1$). This means that, individually, each spin is weakly affected by dissipation. Moreover, we also assume that the phases of the undamped system are not unstable against weak dissipation, i.e., we assume that the nature of the phases do not change by weak dissipation alone.
IV Adiabatic renormalization
This section applies the adiabatic renormalization procedure (leggett_spinboson) to our model system. We remark that it is applicable to the regimes of Ohmic and super-Ohmic dissipation $\left(s\geq 1\right)$. Although not applicable to the sub-Ohmic regime ($s<1$), insights can be gained.
The idea is to integrate out fast oscillators which can adiabatically follow spins tunneling between the up and down $z$-states due to $S^{x}$ operators in the system Hamiltonian (4). Precisely, we integrate out the oscillators within the frequency range $\omega_{l}<\omega_{k,i}<\omega_{c}$. The lower limit $\omega_{l}$ is defined as $\omega_{l}=pg_{i}=p\epsilon_{h,i}h_{i}$, where $p\gg 1$ is a large real number the exact value of which is not important for our purposes. Technically, $p\gg 1$ ensures that we are integrating out only oscillators with frequencies much higher than the tunneling frequencies $g_{i}$ and $h_{i}$ and, thus, can instantaneously follow to the spin dynamics and, hence, be adiabatically eliminated. The net effect is a renormalization of the tunneling parameters $g_{i}$ and $h_{i}.$ To achieve this, we rewrite the system-bath parts of the Hamiltonian by completing the square and separating out the high-frequency components. The resulting Hamiltonian for the high-frequency oscillators is
$H_{0}=$ $\sum_{\omega_{k,i}>\omega_{l}}\left(\frac{p_{1;k,i}^{2}}{2m_{k,i}}+\frac{m_{k,i}\omega_{k,i}^{2}\left(x_{1;k,i}+a_{0;k,i}\eta_{i}^{z}\right)^{2}}{2}\right)$
$+\sum_{\omega_{k,i}>\omega_{l}}\left(\frac{p_{2;k,i}^{2}}{2m_{k,i}}+\frac{m_{k,i}\omega_{k,i}^{2}\left(x_{2;k,i}+a_{0;k,i}\sigma_{i}^{z}\eta_{i}^{z}\right)^{2}}{2}\right),$ (10)
where an irrelevant constant term has been omitted for simplicity and $a_{0;k,i}=\lambda_{k,i}/m_{k,i}\omega_{k,i}^{2}$ is a constant. Notice that $H_{0}$ is a collection independent harmonic oscillators with offset rest positions depending on $\eta_{i}^{z}$ and the other on $\sigma_{i}^{z}\eta_{i}^{z}$ and, thus, its ground state is fourfold degenerate:
$\begin{array}[]{cc}\left|\Psi_{\pm\pm}\right\rangle=&\left|\pm,\pm\right\rangle\bigotimes_{\omega_{k,i}>\omega_{l}}\left|G_{k,i;\pm}^{(1)}\right\rangle\left|G_{k,i;\pm\pm}^{(2)}\right\rangle,\\ \left|\Psi_{\pm\mp}\right\rangle=&\left|\pm,\mp\right\rangle\bigotimes_{\omega_{k,i}>\omega_{l}}\left|G_{k,i;\pm}^{(1)}\right\rangle\left|G_{k,i;\pm\mp}^{(2)}\right\rangle,\end{array}$ (11)
where we are using the notation $\left|\eta_{i}^{z},\sigma_{i}^{z}\right\rangle$. The oscillators ground states have the form
$\left|G_{k,i;\eta_{i}^{z}}^{(1)}\right\rangle=e^{i\eta_{i}^{z}p_{1;k,i}x_{0;k,i}}\left|0_{k,i}^{(1)}\right\rangle,$ (12)
and
$\left|G_{k,i;\eta_{i}^{z}\sigma_{i}^{z}}^{(2)}\right\rangle=e^{i\eta_{i}^{z}\sigma_{i}^{z}p_{2;k,i}x_{0;k,i}}\left|0_{k,i}^{(2)}\right\rangle,$ (13)
with $\left|0_{k,i}^{(\nu)}\right\rangle$ denoting the ground state of a oscillator with rest position at the origin. The tunneling term in the Hamiltonian (4) is
$H_{1}=-h_{i}\sigma_{i}^{x}\left(1+\eta_{i}^{x}\right)-g_{i}\eta_{i}^{x}.$ (14)
The other terms are all diagonal in the ground state and can simply be incorporated to $H_{0}$ latter on. Treating $H_{1}$ perturbatively up to first order, we get the effective Hamiltonian
$H_{\text{eff}}=-\tilde{h}_{i}\sigma_{i}^{x}\left(1+\eta_{i}^{x}\right)-\tilde{g}_{i}\eta_{i}^{x},$ (15)
with renormalized fields
$\tilde{h}_{i}=h_{i}e^{-\frac{2}{\pi}\intop_{p\epsilon_{h,i}h_{i}}^{\omega_{c}}d\omega\frac{F_{i}(\omega)}{\omega^{2}}}=h_{i}e^{-\alpha_{i}\intop_{p\epsilon_{h,i}h_{i}}^{\omega_{c}}d\omega\frac{\omega^{s-2}}{\omega_{c}^{s-1}}},$ (16)
and
$\tilde{g}_{i}=g_{i}e^{-\frac{4}{\pi}\intop_{pg_{i}}^{\omega_{c}}d\omega\frac{F_{i}(\omega)}{\omega^{2}}}=g_{i}e^{-2\alpha_{i}\intop_{pg_{i}}^{\omega_{c}}d\omega\frac{\omega^{s-2}}{\omega_{c}^{s-1}}}.$ (17)
Notice that the effective dissipation parameter for the product variables (related to $g_{i}$) is twice that of the bare spin variables. This may be expected since a spin flip in the product variable $\eta_{i}^{x}=S_{1,i}^{x}S_{2,i}^{x}$ involves changing the rest position of twice the number of oscillators when compared to $\sigma_{i}^{x}=S_{1,i}^{x}$.
Because the effective Hamiltonian (15) keeps its original form, this procedure can be iterated until convergence $g_{i}^{*}$ and $h_{i}^{*}$. As shown in Ref. (leggett_spinboson), in the Ohmic regime ($s=1$), $h_{i}^{*}$ and $g_{i}^{*}$ converge to finite values $h_{i}(p\epsilon_{h}h_{i}/\omega_{c})^{\alpha/\left(1-\alpha\right)}$ and $g_{i}(pg_{i}/\omega_{c})^{2\alpha/\left(1-2\alpha\right)}$ if $\alpha_{i}<1$ and <1/2<1/2, respectively, otherwise, they converge to vanishing values, and tunneling ceases. For super-Ohmic dissipation ($s>1$), $g_{i}^{*}$ and $h_{i}^{*}$ are always finite.
As we are interested in the case of weak dissipation, the bare value of the dissipation parameter $\alpha_{i}\ll 1$. Therefore, dissipation has little effect individually as $h_{i}^{*}\approx h_{i}$ and $g_{i}^{*}\approx g_{i}$. However, near a phase transition, spins at different sites become correlated and, as shown in Ref. (smeared_hoyos_1), the effective dissipation parameter can change under RG. The central question then becomes the following: under what conditions does the dissipation parameter flow to larger values? As we show in the following, this happens only when FM-like RRs appear, and in the FM phase. Surprisingly, not when PROD-like RRs are formed.
V Strong-disorder renormalization group
We now present the SDRG decimation procedure relevant for the $\epsilon_{I}>1$ regime in the presence of dissipation. We follow the methods of the dissipationless case in Ref. (2014_ashkinteller) and include dissipation using the strategy of Ref. (smeared_hoyos_2) which is the following. The renormalization-group decimations are of two types: global and local, and are performed simultaneously when lowering the cutoff energy scale of the system. The global decimation concerns the integration of the fastest bath oscillators and is related to the adiabatic renormalization (see Sec. IV). The local decimation concerns the coarse-grainning of a high-energy localized mode in a disordered system and, thus, is identical to the dissipationless case.
Let $\Omega=\mathrm{max}\left\{K_{i},g_{i},J_{i},h_{i},\omega_{c}/p\right\}$ be the largest energy scale in the renormalized system. The decimation procedure now unfolds in the following steps:
(i) In lowering the energy scale from $\Omega$ to $\Omega-d\Omega$, we integrate out all the oscillators in the interval $\omega_{c}-pd\Omega<\omega_{k,i}<\omega_{c}=p\Omega$. As a result, all tunneling frequencies renormalize to [see Eqs. (16) and (17)]
$\tilde{h}_{i}$ $=h_{i}\left[1-\alpha_{i}\left(\frac{\Omega}{\Omega_{I}}\right)^{s-1}\frac{d\Omega}{\Omega}\right],$ (18)
and
$\tilde{g}_{i}=g_{i}\left[1-2\alpha_{i}\left(\frac{\Omega}{\Omega_{I}}\right)^{s-1}\frac{d\Omega}{\Omega}\right],$ (19)
where $\Omega_{I}$ is the bare $\Omega$. Notice that this is the global decimation procedure mentioned above. In contrast, the remaining steps are local decimations and are identical to the dissipationless case, except when it comes to renormalize the dissipation parameter $\alpha_{i}$.
(ii) Let the two largest energy scales be $g_{i}$ and $K_{i}$. Thus, the unperturbed Hamiltonian is $H_{0}=-K_{i}\sigma_{i}^{z}\sigma_{i+1}^{z}-g{}_{i}\eta_{i}^{x}$ and all other terms in the Hamiltonian (4) containing $\boldsymbol{\sigma}_{i}$, $\boldsymbol{\sigma}_{i+1}$, and $\boldsymbol{\eta}_{i}$ are treated perturbatively up to second order. As a result, the new effective Hamiltonian has the same form as the Hamiltonian (4) but with one fewer site and two fewer spin-1/2 degrees of freedom. The degrees of freedom $\boldsymbol{\sigma}_{i}$ and $\boldsymbol{\sigma}_{i+1}$ are clustered together and replaced by an effective spin-1/2 degree of freedom $\tilde{\boldsymbol{\sigma}}$ upon which the effective field is $\tilde{h}$. In addition, $\boldsymbol{\eta}_{i}$ becomes locked in the $x$-direction and does not contribute to the $z$-magnetization. The effective coupling between the neighboring $\boldsymbol{\eta}$’s is $\tilde{J}$. This is sketched in Fig. 2(a), and the renormalized couplings are
$\tilde{J}=\frac{2J_{i-1}J_{i}}{\tilde{g}_{i}},\quad\tilde{h}=\frac{2h_{i}h_{i+1}}{K_{i}}.$ (20)
Now, we need to compute how the dissipation strength $\alpha$ changes under this decimation. Recalling that
$H_{\text{SB}}=$ $-\eta_{i}^{z}\sum_{k}\lambda_{k,i}x_{1;k,i}-\sigma_{i}^{z}\eta_{i}^{z}\sum_{k}\lambda_{k,i}x_{2;k,i}$
$-\sigma_{i+1}^{z}\eta_{i+1}^{z}\sum_{k}\lambda_{k,i+1}x_{2;k,i+1},$ (21)
and that the doubly degenerated ground state is $\left\{\left|\rightarrow_{i};\uparrow_{i}\uparrow_{i+1}\right\rangle,\left|\rightarrow_{i};\downarrow_{i}\downarrow_{i+1}\right\rangle\right\}$. Then, projecting $H_{\text{SB}}$ onto these states, we find that
$\tilde{H}_{\text{SB}}=$ $-\tilde{\sigma}^{z}\eta_{i+1}^{z}\sum_{k}\lambda_{k,i+1}x_{2;k,i+1}.$ (22)
This is simply because $\left\langle\eta_{i}^{z}\right\rangle=\left\langle\eta_{i}^{z}\sigma_{i}^{z}\right\rangle=0$ and $\left\langle\sigma_{i+1}^{z}\right\rangle=\tilde{\sigma}^{z}$. Actually, this could be obtained in another fashion. Since $g_{i}$ is the largest energy scale, it means that the associated degree of freedom is actually already (or nearly) decoupled from the associated bath via successive adiabatic renormalizations (see Sec. IV). Thus, only the last term in (21) needs to be projected onto the ground-state manifold. Importantly, there is no change in the spectral function of the remaining baths, and, therefore, $\alpha_{i+1}$ remains unchanged.
Figure 2: Decimation procedure when the two largest local energy scales in the system are (a) $g_{2}$ and $K_{2}$, (b) $g_{2}$ and $h_{2}$, and (c) $K_{2}$ and $J_{2}$ [in the main text, these correspond to the decimation cases (ii), (iii) and (iv), respectively]. The couplings to the oscillator baths also renormalized and are not illustrated here.
(iii) Let the two largest energy scales be $g_{i}$ and $h_{i}$. In this case, all tunneling frequencies have fully (or nearly) converged to their fixed-point values $g^{*}$ and $h^{*}$ in the adiabatic renormalization. The SDRG procedure is, thus, the same as the dissipationless case. The unperturbed Hamiltonian to be considered is $H_{0}=-h_{i}\sigma_{i}^{x}\left(1+\eta_{i}^{x}\right)-g_{i}\eta_{i}^{x}$. Projecting the spins $\boldsymbol{\sigma}_{i}$ and $\boldsymbol{\eta}_{i}$ onto the ground state of $H_{0}$ means that both spins are going to be decimated. The renormalized system is that illustrated in Fig. 2(b) and the renormalized quantities are
$\tilde{K}=\frac{K_{i-1}K_{i}}{2h_{i}},\quad\tilde{J}=\frac{J_{i-1}J_{i}}{g_{i}+h_{i}}.$ (23)
Since no spin clusters are formed in this step, the projection of $H_{\text{SB}}$ into the ground state results in zero. Hence, $\alpha$ remains unchanged (except, evidently, for those ones decimated).
(iv) The final case to consider is when $K_{i}$ and $J_{i}$ are the largest energy scales. The unperturbed Hamiltonian is $H_{0}=-K_{i}\sigma_{i}^{z}\sigma_{i+1}^{z}-J_{i}\eta_{i}^{z}\eta_{i+1}^{z}-J_{i}\sigma_{i}^{z}\sigma_{i+1}^{z}\eta_{i}^{z}\eta_{i+1}^{z}$. Projecting $\boldsymbol{\sigma}_{i}$, $\boldsymbol{\sigma}_{i+1}$, $\boldsymbol{\eta}_{i}$, and $\boldsymbol{\eta}_{i+1}$ onto the ground state means that those spins are always aligned and, therefore, can be represented by two effective degrees of freedom $\tilde{\boldsymbol{\sigma}}$ and $\tilde{\boldsymbol{\eta}}$. Fusing them into clusters is illustrated in Fig. 2(c) where the renormalized quantities are
$\tilde{g}=\frac{g_{i}g_{i+1}}{2J_{i}},\quad\tilde{h}=\frac{h_{i}h_{i+1}}{K_{i}+J_{i}}.$ (24)
For this step, we need to take a closer look on what happens to the oscillator baths. The system-bath Hamiltonian is
$H_{\text{SB}}=$ $-\eta_{i}^{z}\sum_{k}\lambda_{k,i}x_{1;k,i}-\sigma_{i}^{z}\eta_{i}^{z}\sum_{k}\lambda_{k,i}x_{2;k,i}$
$-\eta_{i+1}^{z}\sum_{k}\lambda_{k,i+1}x_{1;k,i+1}-\sigma_{i+1}^{z}\eta_{i+1}^{z}\sum_{k}\lambda_{k,i+1}x_{2;k,i+1},$ (25)
and needs to be projected onto the 4-fold-degenerate ground state of $H_{0}$, which is $\left\{\left|\eta;\sigma\right\rangle\right\}$ (where $\eta_{j}^{z}\left|\eta;\sigma\right\rangle=\eta\left|\eta;\sigma\right\rangle$, $\sigma_{j}^{z}\left|\eta;\sigma\right\rangle=\sigma\left|\eta;\sigma\right\rangle$, with $j=i,i+1$ and $\eta,\sigma=\pm 1$ representing the effective spins). Since $H_{\text{SB}}$ is diagonal in all four states, the projection can be easily performed, yielding the following effective Hamiltonian
$\tilde{H}_{\text{SB}}=$ $-\tilde{\eta}^{z}\sum_{k,j}\lambda_{k,j}x_{1;k,j}-\tilde{\sigma}^{z}\tilde{\eta}^{z}\sum_{k,j}\lambda_{k,j}x_{2;k,j},$ (26)
where the sum in the index $j$ run from $i$ to $i+1$. Consequently, the spectral density of the local renormalized bath is
$\tilde{F}(\omega)=\frac{\pi}{2}\sum_{k,j}\frac{\lambda_{k,j}^{2}}{m_{k,j}\omega_{k,j}}\delta(\omega-\omega_{k,j})=\frac{\pi}{2}\tilde{\alpha}\frac{\omega^{s}}{\omega_{I}^{s-1}}e^{-\omega/\omega_{c}},$ (27)
where we can identify the renormalized dissipation strength as
$\tilde{\alpha}=\alpha_{i}+\alpha_{i+1}.$ (28)
It is thus interesting to notice that $\tilde{\alpha}$, increases under the RG flow only when FM-like clusterings take place, i.e., only when $J_{i}$ is decimated. Moreover, from Eq. (28), it becomes clear that the effective dissipation strength increases as the volume of the FM cluster, just like in the case of a direct transition between the PM and FM phases studied in Refs. (smeared_hoyos_1; smeared_hoyos_2). In addition, it is possible to relate the dissipation strength with the effective magnetic moment as both increases with the cluster volume. It is a convenient relation when analyzing the SDRG flow equations for the distributions of couplings, fields and dissipation parameters (smeared_hoyos_2).
VI Phase diagram
We are now able to describe the phase diagram of our model Hamiltonian (5) sketched in Fig. 1. A phase diagram is obtained after analyzing the corresponding strong-disorder renormalization-group flow of the Hamiltonian parameters. A careful reader may note that the decimation cases (i)–(iv) described in Sec. V do not cover all possible decimations. However, they do cover all decimations necessary to analyze the PM-PROD and PROD-FM transitions as shown in Ref. (2014_ashkinteller). To review the undamped case, we start noticing that the PM-PROD transition happens from the competition between the transverse field $h$ and the 4-spin coupling $K$ [see the first sum in the Hamiltonian (4), the other terms are just bystanders]. The relevant decimations are, thus, those described in cases (ii) and (iii) in the previous Section. Comparing the corresponding renormalized values $\tilde{h}$ and $\tilde{K}$ in Eqs. (20) and (23), the transition line happens when the typical values are $2h_{\text{typ}}=K_{\text{typ}}$. Thus, the transition line in the undamped case is at $\delta\approx\epsilon_{I}/2$. For the PROD-FM transition, the competition is between the 2-spin couplings $J$ and $g$ [see the second sum in the Hamiltonian (4)]. The relevant decimations are thus the cases (ii) and (iv). Following the same reasoning as before, the transition line is at $\delta\approx 2/\epsilon_{I}.$
Let us now analyze the effects of weak Ohmic dissipation $\alpha\ll 1$ and $s=1$ in Eq. (8)] in the PM-PROD transition. Since $\alpha\ll 1$ and does not renormalize, the decimation of fast oscillators has little effect on the renormalized values of $h$’s [see Eq. (18)]. Therefore, Ohmic dissipation is a irrelevant perturbation to the PM-PROD quantum phase transition. Its only effect is a slight change of the critical line. This same conclusion applies to the case of super-Ohmic dissipation ($s>1$) which is qualitatively weaker when compared to Ohmic dissipation (leggett_spinboson). Very likely, the same conclusion may also apply to the case of sub-Ohmic dissipation ($s<1$).
The effects of dissipation is much more dramatic in the PROD-FM transition. As pointed out in Secs. II and V, FM-like RRs are formed around that transition. This is because a local FM coupling $J$ can be one of the strongest energy scale in the system [decimation case (iv) in Sec. V]. When this happens, the effective dissipation strength $\tilde{\alpha}$ increases with the volume of the RR [see Eqs. (27) and (28)]. Thus, a sufficiently large RR (and they can be arbitrarily large in a Griffiths phase (fisher_2)) have an effective dissipation parameter $\tilde{\alpha}>\alpha_{c}=1/2$. As a consequence, these larger RRs will cease tunneling ($\tilde{g}\rightarrow 0$) and undergo their own FM transition when Ohmic dissipation is considered. Thus, true local magnetic order develops. As this transition is not a collective effect of the bulk, the transition is said to be smeared. Any residual interaction between these larger RRs will align them together and the total magnetization will not average out. The Griffiths PROD phase of the undamped system now becomes a inhomogeneous FM phase due to Ohmic dissipation. Upon lowering $\delta$, this inhomogeneous FM phase crosses over to a weakly ordered phase (the former Griffiths FM phase of the undamped system) and, finally, crosses over to the conventional FM phase when further lowering $\delta.$ It is important to notice that this phenomenon is akin to the smearing in the PM-FM transition in the random transverse-field Ising chain studied in Refs. (smeared_hoyos_1; smeared_hoyos_2). Thus, both PM-FM ($\epsilon_{I}<1$) and PROD-FM ($\epsilon_{I}>1$) transitions become smeared in the presence of Ohmic dissipation as sketched in Fig. 1(b). We expect that all these conclusions remain valid for sub-Ohmic dissipation as well. For weak super-Ohmic dissipation, on the other hand, all transitions remain sharp because a RR cannot stop tunneling due to the coupling to the oscillator bath regardless the effective value of $\tilde{\alpha}$.
It is interesting to quantify some key observables in those transitions. For the sharp transitions, the universality class is that of the random transverse-field Ising model which is extensively studied. All critical exponents are universal (i.e., independent of the details of the distributions of the coupling constants and fields) and well known. We refer the reader to Ref. (2014_ashkinteller) for the critical behavior of many observables. For the smearing transitions, many observables were quantified in Refs. (smeared_hoyos_1; smeared_hoyos_2). Here, we only quote the magnetization at the onset of the inhomogeneous FM phase: $\overline{\left\langle\sigma_{i}^{z}\right\rangle}=\overline{\left\langle S_{1,i}^{z}\right\rangle}=\overline{\left\langle S_{2,i}^{z}\right\rangle}\sim w^{1/\alpha},$ where $\overline{\left\langle\cdot\right\rangle}$ denotes the disorder and quantum averages. Here, $\alpha$ is the bare dissipation strength and $w$ is the probability of finding a local coupling constant $J_{i}$ being greater than $\gamma_{i}=\max\left\{h_{i}^{*},g_{i}^{*}\right\}$, where $h^{*}=h\left(h/\Omega_{I}\right)^{\frac{\alpha}{1-\alpha}}\approx h$ and, likewise, $g^{*}\approx g$. Precisely, $w=\int_{0}^{\infty}dJP_{J}(J)\int_{0}^{J}d\gamma P_{\gamma}\left(\gamma\right)$. Notice that $w$ plays the role of distance from transition.
VII Discussion and conclusions
In this work, we investigated the effects of dissipation in the phase transitions of the quantum Ashkin-Teller chain in the presence of quenched disorder. The dissipationless model exhibits three quantum phase transitions (see Fig. 1(a)): a FM-PM transition, a FM-PROD transition, and a PM-PROD transtion. All belonging to the universality class of the random transverse-field Ising chain and displaying off-critical quantum Griffiths phases. Introducing dissipation of the quantum fluctuations, surprisingly, does not smear all of these phase transitions. Only the transitions involving the FM phase are smeared (see Fig. 1(b)).
In the more familiar model of the random-transverse field Ising chain, the low-energy physics is dominated by the so-called rare regions in the Griffiths phases. These are regions which are locally FM and weakly coupled to the bulk. Their dynamics are, thus, slow coherent tunneling between the two lowest energy FM states. Smearing due to dissipation occurs because that slow dynamics is completely halted by Ohmic and sub-Ohmic dissipation. The coherent tunneling requires a coherent changing in all oscillators rest position. When that set of oscillators is sufficiently slow, the rare region stop tunneling and develops its own magnetic order regardless the bulk.
It turns out that the slow tunneling of the rare regions in the PM-PROD phase does not require coherent tunneling of all their constituents. This is because the magnetic order of the PROD phase is a composite one: each color spin is disordered, $\left\langle S_{1}^{z}\right\rangle=\left\langle S_{2}^{z}\right\rangle=0$, while the composite magnetic order finite, $\left\langle S_{1}^{z}S_{2}^{z}\right\rangle\neq 0$. Thus, dissipation do not couple directly to the PROD order parameter and, therefore, the transition remains sharp.
It is interesting to consider whether similar situations appear in other contexts. Composite (quadrupole-like) nematic order appears in diverse contexts such as in frustrated magnets (as, for instance, in the context of order-by-disorder in the $J_{1}$-$J_{2}$ Heisenberg model or in easy-plane pyrochlores) and in iron pnictides. Giving the results here uncovered, dissipation due to the coupling to gapless modes (such as that of a metallic Fermi liquid) may not couple to that composite magnetic order, and, consequently, the fluctuations of the associated nematic rare regions may not be overdamped by dissipation.
It is also interesting to test our result in other models. In Appendix A we consider two possible modifications on the dissipation of the quantum fluctuations: one in which the couplings to the oscillator bath is color-dependent and another in which there is only one oscillator bath for both color spins. In both cases, our result remains: the smearing occurs only to the transitions to the FM phase.
Finally, we argue that our results are not restricted to one spatial dimension. Our SDRG method can be applied to any dimension and the smearing (or lack of) will not depend on the underling coordination number.
Acknowledgements.
This work was supported in part by the Brazilian agencies FAPESP and CNPq. J.A.H. thanks IIT Madras for a visiting position under the IoE program which facilitated the completion of this research work. R.N. acknowledges funding from the Center for Quantum Information Theory in Matter and Space-time, IIT Madras, and from the Department of Science and Technology, Government of India, under Grant No. DST/ICPS/QuST/Theme-3/2019/Q69, as well as support from the Mphasis F1 Foundation via the Centre for Quantum Information, Communication, and Computing (CQuICC).
Appendix A Other dissipative models
In this appendix, we consider two modifications of our model Hamiltonian (5). The first modification (see App. A.1) is to introduce a color-dependent anisotropy in the couplings to the different baths. The second modification is to consider that both color spins at a given site couple to the same oscillator bath.
We anticipate that our previous conclusions do not change: only the transitions to the FM phase become smeared.
A.1 Color-dependent couplings
Here, we consider the same model as in Sec. III but with color dependent coupling constants $\lambda_{k}^{(\nu)}$. In this case, the spectral function will be color dependent $F_{i}^{(\nu)}$, implying a color dependent dissipation strength $\alpha_{i}^{(\nu)}$. The adiabatic renormalization procedure can be generalized, the main difference being that the transverse field $h$ renormalizes differently in each chain. Namely, $h$ unravels into $h^{(1)}$ and $h^{(2)}$ in the effective Hamiltonian
$H_{\text{eff}}=-\tilde{h}_{i}^{(1)}\sigma^{x}-\tilde{h}_{i}^{(2)}\sigma^{x}\eta^{x}-\tilde{g}_{i}\eta^{x},$ (29)
with renormalized tunneling constants given by
$\tilde{h}_{i}^{(1,2)}=h_{i}^{(1,2)}\exp\left(-\alpha_{i}^{(1,2)}\intop_{p^{\prime}h_{i}^{(1,2)}}^{\omega_{c}}d\omega\frac{\omega^{s-2}}{\omega_{c}^{s-1}}\right)$ (30)
and
$\tilde{g}_{i}=g_{i}\exp\left(-\left(\alpha_{i}^{(1)}+\alpha_{i}^{(2)}\right)\intop_{pg_{i}}^{\omega_{c}}d\omega\frac{\omega^{s-2}}{\omega_{c}^{s-1}}\right).$ (31)
Note that we recover the color independent result by setting $\alpha_{i}^{(1)}=\alpha_{i}^{(2)}$.
It is necessary to check how the SDRG recursion relations from Sec. V are modified by the unraveling of $h$. Eqs. (18) and (19) in case (i) follows directly from relations (30) and (31). In case (ii), $h^{(1)}$ and $h^{(2)}$ are renormalized in similar ways, namely Eq. (20) is modified to $\tilde{h}^{(1,2)}=\left(h_{i}^{(1)}+h_{i}^{(2)}\right)h_{i+1}^{(1,2)}/K_{i}$, and the recursion relation for $\tilde{J}$ remains unchanged. The same applies to case (iv), with Eq. (24) modified to $\tilde{h}^{(1,2)}=h_{i}^{(1,2)}h_{i+1}^{(1,2)}/\left(K_{i}+J_{i}\right)$ and $\tilde{g}$ remains unchanged. Finally, in decimation case (iii), Eq. (23) modifies to
$\tilde{K}=\frac{K_{i-1}K_{i}}{h_{i}^{(1)}+h_{i}^{(2)}},\quad\tilde{J}^{(1,2)}=\frac{J_{i-1}J_{i}}{g_{i}+h_{i}^{(1,2)}}.$ (32)
Notice, therefore, that $J$ acquires color-dependence, but not $K$. Despite these modifications, the ground state of this step remains the same as that of Sec. V.
To check how dissipation will affect the phase transitions, we project the coupling Hamiltonian $H_{\text{SB}}$ onto the ground states of each decimation case. For cases (ii) and (iii), the coupling Hamiltonian have the same form as the color-independent case (apart from the color dependence of the $\lambda$’s). Therefore, there is no renormalization of the dissipation strength $\alpha$. For case (iv), proceeding analogously to Sec. VI, we get
$H_{\text{eff}}=\tilde{\eta}^{z}\sum_{k}\tilde{\lambda}_{k}^{(1)}\tilde{x}_{1;k}+\tilde{\sigma}^{z}\tilde{\eta}^{z}\sum_{k}\tilde{\lambda}_{k}^{(2)}\tilde{x}_{2;k},$ (33)
where $\tilde{\lambda}_{k}^{(1,2)}=\lambda_{k,i}^{(1,2)}+\lambda_{k,i+1}^{(1,2)}$. This result implies that the renormalization of the dissipation strength in one color is independent of the other color, and they both renormalize similarly to the color-independent case, with $\tilde{\alpha}^{(1,2)}=\alpha_{i}^{(1,2)}+\alpha_{i+1}^{(1,2)}$. Hence, the color dependence of the coupling constant does not affect our conclusions: only the transitions to the FM phase are smeared due to Ohmic and sub-Ohmic dissipation.
A.2 Single dissipation bath
Here, we investigate whether our results on the model Hamiltonian (5) also applies to another related dissipative model. In (5), each spin is coupled to its own set of oscillator bath. We now consider the case in which different color-spins are coupled to the same set of oscillators. Namely,
$H=$ $H_{\text{S}}+\sum_{i}\eta_{i}^{z}(1+\sigma_{i}^{z})\sum_{k}\lambda_{k,i}x_{k,i}$
$+\sum_{k,i}\left(\frac{p_{k,i}^{2}}{2m_{k,i}}+\frac{m_{k,i}\omega_{k,i}^{2}x_{k,i}^{2}}{2}\right)+\triangle V.$ (34)
The first term $H_{\text{S}}$ is our system model the quantum Ashkin-Teller chain in Eq. (4). The first sum quantifies the coupling of the the two color spins to the same bath oscillator at site $i$. [Recall that $\eta_{i}^{z}(1+\sigma_{i}^{z})=S_{1,i}^{z}+S_{3,i}^{z}$.] The second sum is the total mechanical energy of the free oscillators and, finally,
$\triangle V=\frac{1}{2}\sum_{k,i}m_{k,i}\omega_{k,i}^{2}x_{0;k,i}^{2}\left(1+\sigma_{i}^{z}\right)^{2},$
where $x_{0;k,i}=\lambda_{k,i}/m_{k,i}\omega_{k,i}^{2}$ is the counter-term (caldeiraleggett). It must be added if we wish that the coupling to the bath introduces only dissipation, and not effective fields [which appears when completing the square in the last line of Eq. (34)]. To proceed with the adiabatic renormalization, we set
$H_{0}=$ $\eta_{i}^{z}(1+\sigma_{i}^{z})\sum_{\omega_{k,i}>\omega_{l}}\lambda_{k,i}x_{k,i}+\sum_{\omega_{k,i}>\omega_{l}}\left(\frac{p_{k,i}^{2}}{2m_{k,i}}+\frac{m_{k,i}\omega_{k,i}^{2}x_{k,i}^{2}}{2}\right)$
$+\sum_{\omega_{k,i}>\omega_{l}}\frac{m_{k,i}\omega_{k,i}^{2}x_{0;k,i}^{2}\left(1+\sigma_{i}^{z}\right)^{2}}{2}$
$=$ $\sum_{\omega_{k,i}>\omega_{l}}\left(\frac{p_{k,i}^{2}}{2m_{k,i}}+\frac{m_{k,i}\omega_{k,i}^{2}\left(x_{k,i}+x_{0k,i}\eta_{i}^{z}(1+\sigma_{i}^{z})\right)^{2}}{2}\right),$ (35)
and treat the tunneling terms of the $i$-th site in first order of perturbation theory. The definition of $\omega_{l}$ is the same as in Sec. IV. The 4-fold degenerated ground state of $H_{0}$ is given by
$\left|\Psi_{++}\right\rangle=$ $\left|+,+\right\rangle\prod_{\omega_{k,i}>\omega_{l}}\left|G_{k,i;++}\right\rangle,$
$\left|\Psi_{-+}\right\rangle=$ $\left|-,+\right\rangle\prod_{\omega_{k,i}>\omega_{l}}\left|G_{k,i;-+}\right\rangle,$ (36)
$\left|\Psi_{\pm-}\right\rangle=$ $\left|\pm,-\right\rangle\prod_{\omega_{k,i}>\omega_{l}}\left|0_{k,i}\right\rangle,$
where we are using the same notation as Sec. IV (first and indices to $\eta$ and $\sigma$ variables, respectively). The oscillators’ ground states are
$\left|G_{k,i;\eta_{i}^{z}\sigma_{i}^{z}}\right\rangle=e^{2i\eta_{i}^{z}\sigma_{i}^{z}p_{k,i}x_{0;k,i}}\left|0_{k,i}\right\rangle,$ (37)
and $\left|0_{k,i}\right\rangle$ is the ground state of the oscillator with the rest position at the origin. Projecting the perturbation
$H_{1}=-h_{i}\sigma_{i}^{x}(1+\eta_{i}^{x})-g_{i}\eta_{i}^{x},$ (38)
onto the ground state manyfold (the corresponding matrix element being $H_{\text{eff}\eta^{z}\sigma^{z},\eta^{\prime z}\sigma^{\prime z}}=\left\langle\Psi_{\eta^{z}\sigma^{z}}\right|H_{1}\left|\Psi_{\eta^{\prime z}\sigma^{\prime z}}\right\rangle$), we find the effective Hamiltonian
$H_{\text{eff}}=$ $-\tilde{h_{i}}\sigma_{i}^{x}\left(1+\eta_{i}^{x}\right)-g_{i}\eta_{i}^{x}\left(\frac{1-\sigma_{i}^{z}}{2}\right)$
$-\tilde{g_{i}}\eta_{i}^{x}\left(\frac{1+\sigma_{i}^{z}}{2}\right)$ (39)
with the renormalized tunneling frequencies
$\tilde{h}_{i}=h_{i}\exp\left(-\alpha_{i}\intop_{p^{\prime}h_{i}}^{\omega_{c}}d\omega\frac{\omega^{s-2}}{\omega_{c}^{s-1}}\right),$ (40)
and
$\tilde{g}_{i}=g_{i}\exp\left(-4\alpha_{i}\intop_{pg_{i}}^{\omega_{c}}d\omega\frac{\omega^{s-2}}{\omega_{c}^{s-1}}\right).$ (41)
From Eq. (39), it follows that the tunneling frequency $g_{i}$ splits in two parts: one that decreases under adiabatic renormalization and other that remains unaffected. This occurs because both states $\left|\Psi_{\pm-}\right\rangle$ are centered at the same position, such that the tunneling between states $\left|\Psi_{+-}\right\rangle$ and $\left|\Psi_{--}\right\rangle$, driven by the term $g_{i}\eta_{i}^{x}$ in $H_{1}$, does not change the equilibrium position of the oscillators. As a result, the tunneling frequency $g_{i}$ remains unaffected in this case. However, the $g_{i}\eta_{i}^{x}$ term also promotes tunneling between the states $\left|\Psi_{++}\right\rangle$ and $\left|\Psi_{-+}\right\rangle$, which are centered at different positions. For this tunneling process, the frequency $g_{i}$ undergoes a renormalization, as described in Eq. (41).
To iterate the procedure, we need to generalize the perturbation (38) as
$H_{1}=$ $-h_{i}\sigma_{i}^{x}\left(1+\eta_{i}^{x}\right)-g_{i}^{(a)}\eta_{i}^{x}\left(\frac{1-\sigma_{i}^{z}}{2}\right)$
$-g_{i}^{(b)}\eta_{i}^{x}\left(\frac{1+\sigma_{i}^{z}}{2}\right)$ (42)
and start with $g_{i}^{(a)}=g_{i}^{(b)}$. In this way, the effective Hamiltonian is
$H_{\text{eff}}=$ $-\tilde{h_{i}}\sigma_{i}^{x}\left(1+\eta_{i}^{x}\right)-g_{i}^{(a)}\eta_{i}^{x}\left(\frac{1-\sigma_{i}^{z}}{2}\right)$
$-\tilde{g_{i}}^{(b)}\eta_{i}^{x}\left(\frac{1+\sigma_{i}^{z}}{2}\right),$ (43)
with $\tilde{h}_{i}$ and $\tilde{g}_{i}^{(b)}$ given, respectively, by Eqs. (40) and (41).
For the SDRG, we define the largest energy scale as $\Omega=\mathrm{max}\left\{K_{i},\left(g_{i}^{(a)}+g_{i}^{(b)}\right)/2,J_{i},h_{i},\omega_{c}/p\right\}$, and proceed just as in Sec. V. Decimation case (i) follows directly from relations (40) and (41).
For the decimation case (ii), the unperturbed Hamiltonian is $H_{0}=-K_{i}\sigma_{i}^{z}\sigma_{i+1}^{z}-\frac{\left(g_{i}^{(a)}+g_{i}^{(b)}\right)}{2}\eta_{i}^{x}$. The additional term $\left(g_{i}^{(a)}-g_{i}^{(b)}\right)\eta_{i}^{x}\sigma_{i}^{z}/2$ (initially vanishing) is diagonal in the ground state of $H_{0}$ and contributes only with an unimportant constant, and the recursion relations (20) become
$\tilde{J}=\frac{4J_{i-1}J_{i}}{g_{i}^{(a)}+g_{i}^{(b)}},\quad\tilde{h}=\frac{2h_{i}h_{i+1}}{K_{i}}.$ (44)
For the decimation case (iii), the unperturbed Hamiltonian is $H_{0}=-h_{i}\sigma_{i}^{x}\left(1+\eta_{i}^{x}\right)-\frac{\left(g_{i}^{(a)}+g_{i}^{(b)}\right)}{2}\eta_{i}^{x}$ and, adding the new term to $H_{1}$, the recursions (23) become
$\tilde{K}=\frac{K_{i-1}K_{i}}{2h_{i}},\quad\tilde{J}=\frac{2J_{i-1}J_{i}}{2h_{i}+g_{i}^{(a)}+g_{i}^{(b)}}.$ (45)
For the decimation step (iv), nothing changes. The recursion relations (24) remain but with a trivial modification of $g^{(a,b)}$.
Now, we can analyze the impact of dissipation on the phase transitions. Since the steps of the SDRG process remain similar to those outlined in Sec. V, the projection of
$H_{\text{SB}}=\eta_{i}^{z}(1+\sigma_{i}^{z})\sum_{k}\lambda_{k,i}x_{k,i}+\eta_{i+1}^{z}(1+\sigma_{i+1}^{z})\sum_{k}\lambda_{k,i+1}x_{k,i+1}$ (46)
onto the ground states corresponding to decimation cases (ii) and (iii) does not alter the dissipation strength $\alpha$. In contrast, for the case (iv), the coupling Hamiltonian becomes
$\tilde{H}_{\text{SB}}=\tilde{\eta}^{z}(1+\tilde{\sigma}^{z})\sum_{k}\left(\lambda_{k,i}+\lambda_{k,i+1}\right)\tilde{x}_{k},$ (47)
which leads to the same renormalization of $\lambda_{k,i}$ (and, consequently, $\alpha_{i}$) just like in the situation analyzed in Sec. V. Thus, the effect of dissipation on the phase transitions mirrors the earlier conclusions: the FM-PROD transition is smeared, while the PM-PROD transition remains sharp.

### v2

@@
 This is due to a cancellation of the dissipation effects on the nontrivial nature of the intertwined order parameter of one of the phases.
+Published in Phys. Rev. B 113, 125128 (2026) DOI: 10.1103/mnwg-hd8m
 I Introduction
-Quenched disorder underlies a manifold of profound effects near quantum phase transitions: For instance, disorder-induced inhomogeneities lead to exotic universality classes such as those typified by the infinite randomness critical point (fisher_1; fisher_2), and the strong off-critical Griffiths singularities (griffiths_original; mccoy-prl69).
-Furthermore, in certain situations these inhomogeneities could lead to the destruction or smearing (smeared_hoyos_1; smeared_hoyos_2) of the putative quantum phase transition.
+Quenched disorder underlies a manifold of profound effects near quantum phase transitions: For instance, disorder-induced inhomogeneities lead to exotic universality classes such as those typified by the infinite randomness critical point [1, 2], and the strong off-critical Griffiths singularities [3, 4].
+Furthermore, in certain situations these inhomogeneities could lead to the smearing [5, 6] of the sharp quantum phase transition.
 The moorings of the effects described above is centered around the so-called rare regions (RRs): arbitrarily large regions in space which, due to a statistically rare configuration of the quenched-disordered couplings, is locally ordered and weakly coupled to the bulk.
-These RRs, characterized by slow dynamics, dominate the low-energy behavior of physical observables (glass_1; glass_2), (for reviews, see Refs.
-(vojta2006rare; igloi-review; vojta2013phases; igloi-monthus-review2)).
+These RRs, characterized by slow dynamics, dominate the low-energy behavior of physical observables [7, 8], (for reviews, see Refs.
+[9, 10, 11, 12]).
 In practice it is important to realize that real materials are seldom described as closed systems.
-Thus, it is of paramount importance to understand how the phases and phase transitions gets affected when the system is connected to a dissipative reservoir.
-The combined effects of quenched disorder and dissipation have been investigated in the Griffiths phase where it was established that sufficiently strong dissipation can completely damp the fluctuations of a RR (millis-morr-schmalian-prl01; vojta-prl03; vojta-schmalian-prb05; vojta-hoyos-prl14).
-Further, Strong-Disorder Renormalization-Group (SDRG) based analytical study established that for the case of the Random Transverse-Field Ising Model (RTFIM), in any dimension, the critical point is entirely destroyed by smearing when dissipation is Ohmic or Subohmic (smeared_hoyos_1; smeared_hoyos_2).
-Physically, the smearing phenomena occurs because dissipation freezes the dynamics of sufficiently large RRs, independent of the bulk, i.e., each rare region orders independently from the bulk, leading to local phase transitions that can take place at different values of the control parameter.
-Thus, the paramagnetic Griffiths phase becomes an inhomogeneously ordered ferromagnetic phase.
-Notably, experimental evidence of smeared phase transitions has been observed in the compounds Sr1-xCaxRuO3 (experimental_1) and CePd1-xRhx (experimental_2).
-More recently, smeared phase transitions have also been identified in a numerical study of strange metals (strangemetals), and have also potentially seen in low-dimensional superconductors (kaur-etal-njp24).
+Thus it is of paramount importance to understand how the phases and phase transitions get affected when the system is connected to a dissipative reservoir.
+Initially, it was realized that sufficiently strong dissipation can completely damp the fluctuations of a RR [13, 14, 15, 16], thus establishing that the combined effect of dissipation and disorder is nonperturbative and could round the sharp phase transition by smearing.
+To quantify the resulting phase and phase transition, interactions among the RRs should be taken into account and it was shown that this could be accomplished via a the strong-disorder renormalization-group (SDRG) method [17, 18].
+Later, it was established that for the case of the random transverse-field Ising model (RTFIM), in any dimension, the critical point is entirely destroyed by smearing when dissipation is Ohmic or Sub-Ohmic [5, 6].
+Physically, the smearing phenomenon occurs because dissipation freezes the dynamics of sufficiently large RRs, independent of the bulk, i.e., each rare region orders independently from the bulk, leading to local phase transitions that can take place at different values of the control parameter.
+Thus the paramagnetic Griffiths phase becomes an inhomogeneously ordered ferromagnetic phase.
+Many other numerical studies (either SDRG or Monte Carlo) have confirmed this conclusion [11, 12].
+Notably, experimental evidence of smeared phase transitions has been observed in the compounds Sr1-xCaxRuO3 [19] and CePd1-xRhx [20].
+More recently, smeared phase transitions have also been identified in a numerical study of strange metals [21], and have also potentially been seen in low-dimensional superconductors [22].
 This work belongs to the genre of problems that we have discussed in the previous paragraph: Namely, we study the dissipative random quantum Ashkin-Teller (RQAT) model.
-Besides the usual paramagnetic (PM) and ferromagnetic (FM) phase, this model features an additional phase, called the product phase, which exhibits a composite (or intertwined) order (1981_ashkinteller).
+Besides the usual paramagnetic (PM) and ferromagnetic (FM) phases, this model features an additional phase, called the product phase, which exhibits a composite (or intertwined) order [23].
 We inquire whether the transitions between this product phase and the other phases are smeared by the combined effects of dissipation and disorder.
@@
 Surprisingly, even though dissipation is coupled to the original order parameter fields, it does not couple generically to the intertwined order parameter field and, therefore, does not damp the quantum fluctuations of the associated rare regions that form near the transition line between the product and PM phases.
-Finally, besides this fundamental motivation, it is worth mentioning that the Ashkin-Teller model has applications in other contexts such as to describe layers of atoms absorbed on surfaces (bak-etal-prl85), current loops in high-Tc superconductors (aji-varma-prl07), and the elastic response of DNA molecules (chang-wang-zheng-ctp08).
+Finally, besides this fundamental motivation, it is worth mentioning that the Ashkin-Teller model has applications in other contexts such as to describe layers of atoms absorbed on surfaces [24], current loops in high-$T_{c}$ superconductors [25], and the elastic response of DNA molecules [26].
 The understanding of dissipation in these contexts is another motivation for our work.
@@
 II, introduce the model studied in Sec.
-III, apply the adiabatic renormalization-group and the strong-disorder renormalization-group methods to our problem in Secs.
+III, and apply the adiabatic renormalization-group and the strong-disorder renormalization-group methods to our problem in Secs.
 IV and V, respectively.
@@
 VII.
-Other possible models of dissipation are considered in Appendix A and we find that our conclusions remain qualitatively correct.
-II Review of the disspationless model
-The Hamiltonian of the one-dimensional random quantum Ashkin-Teller (RQAT) model is (1943_ashkinteller; 1981_ashkinteller; 2001_ashkinteller; 2008_ashkinteller; 2014_ashkinteller)
+Other possible models of dissipation are considered in the Appendix A and we find that our conclusions remain qualitatively correct.
+II Review of the dissipationless model
+II.1 Original variables
+The Hamiltonian of the one-dimensional random quantum Ashkin-Teller (RQAT) model is [27, 23, 28, 29, 30]
 $H_{\text{S}}=$ $-\sum_{\nu=1}^{2}\sum_{i}\left(J_{i}S_{\nu,i}^{z}S_{\nu,i+1}^{z}+h_{i}S_{\nu,i}^{x}\right)$
@@
 The coupling strengths are characterized by the ratios $\epsilon_{h,i}=g_{i}/h_{i}$ and $\epsilon_{J,i}=K_{i}/J_{i}$.
-For simplicity, we assume that the (bare) values $\epsilon_{h,i}=\epsilon_{J,i}=\epsilon_{I}>0$, i.e., they are uniform and site independent.
-However, they flow under the renormalization-group transformations and become random variables (2014_ashkinteller).
+For simplicity, we assume that the (bare) values $\epsilon_{h,i}=\epsilon_{J,i}=\epsilon_{I}>0$; i.e., they are uniform and site independent.
+However, they flow under the renormalization-group transformations and become random variables [30].
 The low-energy physical behavior of the model (1) can be divided into two regimes: the weak-coupling regime ($\epsilon_{I}<1$) and the strong-coupling regime ($\epsilon_{I}>1$).
-In the low-energy limit of the weak-coupling regime, the two Ising chains effectively decouple (2001_ashkinteller; 2008_ashkinteller) and the resulting physics is that of two identical RTFI chains.
+In the low-energy limit of the weak-coupling regime, the two Ising chains effectively decouple [28, 29] and the resulting physics is that of two identical RTFI chains.
 Therefore, there is a direct quantum phase transition between a ferromagnet (where the quantum average $\left\langle S_{\nu,i}^{z}\right\rangle\neq 0$) and paramagnet (where $\left\langle S_{\nu,i}^{z}\right\rangle=0$).
 The transition, which occurs when $\prod J_{i}=\prod h_{i}$, is of infinite-randomness type and is surrounded by the Griffiths ferro- and paramagnetic phases.
-In the strong-coupling regime $\epsilon_{I}>1$, on the other hand, this transition is through an intervening product phase exhibiting a composite order, i.e., $\left\langle S_{\nu,i}^{z}\right\rangle=0$, but $\left\langle S_{1,i}^{z}S_{2,i}^{z}\right\rangle\neq 0$.
+In the strong-coupling regime $\epsilon_{I}>1$, on the other hand, this transition is through an intervening product phase exhibiting a composite order; i.e., $\left\langle S_{\nu,i}^{z}\right\rangle=0$, but $\left\langle S_{1,i}^{z}S_{2,i}^{z}\right\rangle\neq 0$.
 The transition between the ferromagnet and the product phase is also of infinite-randomness type surrounded by quantum Griffiths ferromagnetic and product phases.
 By duality, [^1] The Hamiltonian (1) is invariant under the duality transformation: $S_{\nu,i}^{z}S_{\nu,i+1}^{z}\rightarrow\tau_{\nu,i}^{x}$, $S_{\nu,i}^{x}\rightarrow\tau_{\nu,i}^{z}\tau_{\nu,i+1}^{z}$ (here, $\tau^{x}$ and $\tau^{z}$ are the dual Pauli operators), $J_{i}\leftrightarrow h_{i}$, and $K_{i}\leftrightarrow g_{i}$.
-the transition between the paramagnetic and product phases is in the same universality class (2014_ashkinteller).
+the transition between the paramagnetic and product phases is in the same universality class [30].
+II.2 Product variable
 This result is better illustrated by reformulating the problem in terms of “product” variables $\boldsymbol{\sigma}_{i}$ (which are also spin-1/2 operators represented by Pauli matrices) and one of the “original” spin variable $\boldsymbol{\eta}_{i}$, defined as
@@
 $-\sum_{i}\left(J_{i}\sigma_{i}^{z}\sigma_{i+1}^{z}\eta_{i}^{z}\eta_{i+1}^{z}+h_{i}\sigma_{i}^{x}\eta_{i}^{x}\right).$ (4)
-From the first and second sums of the Hamiltonian (4), it is clear that, when $\epsilon_{I}\gg 1$ and sufficiently near to the self-duality line $\left(\delta\approx 0\right)$, the system is in a phase with long-range order in the $\boldsymbol{\sigma}$ variables ($\left\langle\sigma^{z}\right\rangle\neq 0$) and disordered in the $\boldsymbol{\eta}$ variables ($\left\langle\eta^{z}\right\rangle=0$).
-This implies that long-range order develops in the product variables $\left\langle S_{1}^{z}S_{2}^{z}\right\rangle\neq 0$, while the spins remain disordered in each individual color $\left\langle S_{1}^{z}\right\rangle=\left\langle S_{2}^{z}\right\rangle=0$ (1943_ashkinteller; 1981_ashkinteller).
-Thus, the moniker product (PROD) phase.
+From the first and second sums of the Hamiltonian (4), it is clear that, when $\epsilon_{I}\gg 1$ and sufficiently near to the self-duality line $\left(\delta\equiv\left\langle\ln\left(h/J\right)\right\rangle\approx 0\right)$, the system is in a phase with long-range order in the $\boldsymbol{\sigma}$ variables ($\left\langle\sigma^{z}\right\rangle\neq 0$) and disordered in the $\boldsymbol{\eta}$ variables ($\left\langle\eta^{z}\right\rangle=0$).
+This implies that long-range order develops in the product variables $\left\langle S_{1}^{z}S_{2}^{z}\right\rangle\neq 0$, while the spins remain disordered in each individual color $\left\langle S_{1}^{z}\right\rangle=\left\langle S_{2}^{z}\right\rangle=0$ [27, 23]- thus, the moniker “product (PROD) phase”.
 A caveat is needed here.
 The terminology $\boldsymbol{\sigma}_{i}$ and $\boldsymbol{\eta}_{i}$ as a product and original variables is due to the lack of a better one.
-It is precise if one considers only the $z$-component of the original spins [see Eq.
+It is precise if one considers only the $z$ component of the original spins [see Eq.
 (2)].
-Regarding the $x$-component [see Eq.
+Regarding the $x$ component [see Eq.
 (3)], however, the terminology is the other way round.
-Regarding the $y$-component, both could be regarded as product variables.
+Regarding the $y$ component, both could be regarded as product variables.
 In sum, both $\boldsymbol{\sigma}_{i}$ and $\boldsymbol{\eta}_{i}$ are nontrivial compositions of the original spin variables.
-We stick with this terminology because the $z$-component is the one defining the order parameter field of the transition.
-In the presence of quenched disorder, it was found that all quantum phase transitions belong to the same universality class (that of the random transverse-field Ising chain) and are surrounded by the associated quantum Griffiths phases (2014_ashkinteller) as illustrated in the phase diagram Fig.
+We stick with this terminology because the $z$ component is the one defining the order parameter field of the transition.
+II.3 Phase diagram: Dissipationless case
+In the presence of quenched disorder, it was found that all quantum phase transitions belong to the same universality class (that of the random transverse-field Ising chain) and are surrounded by the associated quantum Griffiths phases [30] as illustrated in the phase diagram Fig.
 1(a).
-Figure 1: (a) Schematic phase diagram of the RQAT chain without dissipation.
-The dashed lines indicate crossovers from conventional FM, PM and product phases to Griffiths phases.
+Figure 1: (a) Schematic phase diagram of the RQAT chain without dissipation ($\alpha=0$, discussed in Sec.
+II).
 GFM and GPM denote the Griffiths ferromagnetic and paramagnetic phases, respectively.
@@
 GPROD denotes the Griffiths product phase, where rare regions form in the $\sigma$ variable near the PM transition and in the $\eta$ variable near the FM transition.
-(b) Schematic phase diagram of the RQAT chain with dissipation.
+(b) Schematic phase diagram of the RQAT chain with dissipation ($\alpha\neq 0$, discussed in Sec.
+VI).
 The transitions where rare regions are formed in the original spin variable are smeared.
 In the strong-coupling regime, the GFM phase is replaced by a weakly ordered (WO) phase and the GPROD phase is replaced by an inhomogeneous ferromagnetic (IFM) phase.
+In all cases, thick solid lines represent second-order quantum phase transitions.
+The dot-dashed line in (b) represents the smeared transition between the PROD and FM phases.
+The dashed lines indicate crossovers, not transitions.
+Different colors represent different phases.
 It is important to distinguish between the two sorts of Griffiths phases in the region $\epsilon_{I}>1$.
-Surrounding the PM–PROD transition, the Griffiths singularities in both the Griffiths PM and Griffiths PROD phases are due to rare regions locally in a PROD phase and weakly interacting with the bulk.
-In RRs of this kind, the $z$-component of the original spins fluctuate rapidly and independently of the nearby spins.
+Surrounding the PM-PROD transition, the Griffiths singularities in both the Griffiths PM and Griffiths PROD phases are due to rare regions locally in a PROD phase and weakly interacting with the bulk.
+In RRs of this kind, the $z$ components of the original spins fluctuate rapidly and independently of the nearby spins.
 However, they fluctuate together with the other color spin in the same site.
-On the other hand, surrounding the PROD-FM transition, the low-energy behavior of both the Griffiths PROD and FM phases are dominated by RRs locally in the FM phase and weakly coupled to the bulk.
-Near the multicritical point, both RRs coexist in the region $\epsilon_{I}>1$ and, therefore, give rise to the double Griffiths (DG) phase (2014_ashkinteller).
-It is worth noting that this DG phase have been seen in a density-matrix renormalization-group study (chatelain-voliotis-epjb16).
+On the other hand, surrounding the PROD-FM transition, the low-energy behavior of both the Griffiths PROD and FM phases is dominated by RRs locally in the FM phase and weakly coupled to the bulk.
+Near the multicritical point, both RRs coexist in the region $\epsilon_{I}>1$ and, therefore, give rise to the double Griffiths (DG) phase [30].
+It is worth noting that this DG phase has been seen in a density-matrix renormalization-group study [31].
 Evidently, these two phase transitions and the associated Griffiths phases are related by duality.
@@
 As a result, only the FM-like RRs will be dramatically affected by dissipation yielding to the smearing of the transition as sketched in Fig.
-c.
-III A simple dissipative model
+1(b) (see further discussions in Sec.
+VI).
+III Simple dissipative model
 We introduce dissipation by coupling each spin to its own set of identical bosonic baths. The Hamiltonian thus becomes
 $H=H_{\text{S}}+\sum_{\nu=1}^{2}H_{\text{B}}^{(\nu)}+\sum_{\nu=1}^{2}H_{\text{SB}}^{(\nu)}$ (5)
-where $H_{\text{S}}$ is the RQAT Hamiltonian defined in Eq. (1). The bath component are independent harmonic oscillators, and thus,
+where $H_{\text{S}}$ is the RQAT Hamiltonian defined in Eq. (1). The bath components are independent harmonic oscillators and, thus,
 $H_{\text{B}}^{(\nu)}=\sum_{k,i}\left(\frac{p_{\nu;k,i}^{2}}{2m_{k,i}}+\frac{m_{k,i}\omega_{k,i}^{2}x_{\nu;k,i}^{2}}{2}\right).$ (6)
-Here, the oscillator frequencies $\omega_{k,i}$ and masses $m_{k,i}$ are assumed to be color-independent.
-Finally, following the Caldeira-Leggett model (caldeiraleggett; leggett_spinboson; smeared_hoyos_1; smeared_hoyos_2), the system-bath coupling is
+Here, the oscillator frequencies $\omega_{k,i}$ and masses $m_{k,i}$ are assumed to be color independent. Finally, following the Caldeira-Leggett model [32, 33, 5, 6], the system-bath coupling is
 $H_{\text{SB}}^{(\nu)}=\sum_{i}S_{\nu,i}^{z}\sum_{k}\lambda_{k,i}x_{\nu;k,i},$ (7)
-where the constant $\lambda_{k,i}$ controls the strength of the coupling. For the moment, we consider color-independent coupling constants. Later, we discuss the effects of color-dependence.
+where the constant $\lambda_{k,i}$ controls the strength of the coupling. For the moment, we consider color independent coupling constants. Later, we discuss the effects of color dependence.
 The relevant properties of the baths are described by their spectral densities $F_{i}(\omega)$, defined as
@@
 In what follows, we follow Refs.
-(smeared_hoyos_1; smeared_hoyos_2) to generalize the adiabatic renormalization procedure (leggett_spinboson) to the present case.
+[5, 6] to generalize the adiabatic renormalization procedure [33] to the present case.
 Importantly, we assume weak dissipation ($\alpha_{i}\ll 1$).
 This means that, individually, each spin is weakly affected by dissipation.
-Moreover, we also assume that the phases of the undamped system are not unstable against weak dissipation, i.e., we assume that the nature of the phases do not change by weak dissipation alone.
+Moreover, we also assume that the phases of the undamped system are not unstable against weak dissipation, i.e., we assume that the nature of the phases does not change by weak dissipation alone.
 IV Adiabatic renormalization
-This section applies the adiabatic renormalization procedure (leggett_spinboson) to our model system.
+This section applies the adiabatic renormalization procedure [33] to our model system.
 We remark that it is applicable to the regimes of Ohmic and super-Ohmic dissipation $\left(s\geq 1\right)$.
@@
 where an irrelevant constant term has been omitted for simplicity and $a_{0;k,i}=\lambda_{k,i}/m_{k,i}\omega_{k,i}^{2}$ is a constant.
-Notice that $H_{0}$ is a collection independent harmonic oscillators with offset rest positions depending on $\eta_{i}^{z}$ and the other on $\sigma_{i}^{z}\eta_{i}^{z}$ and, thus, its ground state is fourfold degenerate:
-$\begin{array}[]{cc}\left|\Psi_{\pm\pm}\right\rangle=&\left|\pm,\pm\right\rangle\bigotimes_{\omega_{k,i}>\omega_{l}}\left|G_{k,i;\pm}^{(1)}\right\rangle\left|G_{k,i;\pm\pm}^{(2)}\right\rangle,\\ \left|\Psi_{\pm\mp}\right\rangle=&\left|\pm,\mp\right\rangle\bigotimes_{\omega_{k,i}>\omega_{l}}\left|G_{k,i;\pm}^{(1)}\right\rangle\left|G_{k,i;\pm\mp}^{(2)}\right\rangle,\end{array}$ (11)
-where we are using the notation $\left|\eta_{i}^{z},\sigma_{i}^{z}\right\rangle$. The oscillators ground states have the form
+Notice that $H_{0}$ is a collection of independent harmonic oscillators with offset rest positions depending on $\eta_{i}^{z}$ and the other on $\sigma_{i}^{z}\eta_{i}^{z}$ and, thus, its ground state is fourfold degenerate:
+$\left|\Psi_{\pm\pm}\right\rangle$ $=$ $\left|\pm,\pm\right\rangle\bigotimes_{\omega_{k,i}>\omega_{l}}\left|G_{k,i;\pm}^{(1)}\right\rangle\left|G_{k,i;\pm\pm}^{(2)}\right\rangle,$ (11)
+$\left|\Psi_{\pm\mp}\right\rangle$ $=$ $\left|\pm,\mp\right\rangle\bigotimes_{\omega_{k,i}>\omega_{l}}\left|G_{k,i;\pm}^{(1)}\right\rangle\left|G_{k,i;\pm\mp}^{(2)}\right\rangle,$ (12)
+where we are using the notation $\left|\eta_{i}^{z},\sigma_{i}^{z}\right\rangle$. The oscillators’ ground states have the form
 $\left|G_{k,i;\eta_{i}^{z}}^{(1)}\right\rangle=e^{i\eta_{i}^{z}p_{1;k,i}x_{0;k,i}}\left|0_{k,i}^{(1)}\right\rangle,$ (13)
@@
 $\left|G_{k,i;\eta_{i}^{z}\sigma_{i}^{z}}^{(2)}\right\rangle=e^{i\eta_{i}^{z}\sigma_{i}^{z}p_{2;k,i}x_{0;k,i}}\left|0_{k,i}^{(2)}\right\rangle,$ (14)
-with $\left|0_{k,i}^{(\nu)}\right\rangle$ denoting the ground state of a oscillator with rest position at the origin. The tunneling term in the Hamiltonian (4) is
+with $\left|0_{k,i}^{(\nu)}\right\rangle$ denoting the ground state of an oscillator with rest position at the origin. The tunneling term in the Hamiltonian (4) is
 $H_{1}=-h_{i}\sigma_{i}^{x}\left(1+\eta_{i}^{x}\right)-g_{i}\eta_{i}^{x}.$ (15)
-The other terms are all diagonal in the ground state and can simply be incorporated to $H_{0}$ latter on. Treating $H_{1}$ perturbatively up to first order, we get the effective Hamiltonian
+The other terms are all diagonal in the ground state and can simply be incorporated to $H_{0}$ later on. Treating $H_{1}$ perturbatively up to first order, we get the effective Hamiltonian
 $H_{\text{eff}}=-\tilde{h}_{i}\sigma_{i}^{x}\left(1+\eta_{i}^{x}\right)-\tilde{g}_{i}\eta_{i}^{x},$ (16)
 with renormalized fields
-$\tilde{h}_{i}=h_{i}e^{-\frac{2}{\pi}\intop_{p\epsilon_{h,i}h_{i}}^{\omega_{c}}d\omega\frac{F_{i}(\omega)}{\omega^{2}}}=h_{i}e^{-\alpha_{i}\intop_{p\epsilon_{h,i}h_{i}}^{\omega_{c}}d\omega\frac{\omega^{s-2}}{\omega_{c}^{s-1}}},$ (16)
+$\tilde{h}_{i}=h_{i}e^{-\frac{2}{\pi}\intop_{p\epsilon_{h,i}h_{i}}^{\omega_{c}}d\omega\frac{F_{i}(\omega)}{\omega^{2}}}\approx h_{i}e^{-\alpha_{i}\intop_{p\epsilon_{h,i}h_{i}}^{\omega_{c}}d\omega\frac{\omega^{s-2}}{\omega_{c}^{s-1}}},$ (17)
 and
-$\tilde{g}_{i}=g_{i}e^{-\frac{4}{\pi}\intop_{pg_{i}}^{\omega_{c}}d\omega\frac{F_{i}(\omega)}{\omega^{2}}}=g_{i}e^{-2\alpha_{i}\intop_{pg_{i}}^{\omega_{c}}d\omega\frac{\omega^{s-2}}{\omega_{c}^{s-1}}}.$ (17)
+$\tilde{g}_{i}=g_{i}e^{-\frac{4}{\pi}\intop_{pg_{i}}^{\omega_{c}}d\omega\frac{F_{i}(\omega)}{\omega^{2}}}\approx g_{i}e^{-2\alpha_{i}\intop_{pg_{i}}^{\omega_{c}}d\omega\frac{\omega^{s-2}}{\omega_{c}^{s-1}}}.$ (18)
 Notice that the effective dissipation parameter for the product variables (related to $g_{i}$) is twice that of the bare spin variables.
 This may be expected since a spin flip in the product variable $\eta_{i}^{x}=S_{1,i}^{x}S_{2,i}^{x}$ involves changing the rest position of twice the number of oscillators when compared to $\sigma_{i}^{x}=S_{1,i}^{x}$.
-Because the effective Hamiltonian (15) keeps its original form, this procedure can be iterated until convergence $g_{i}^{*}$ and $h_{i}^{*}$.
+Because the effective Hamiltonian (16) keeps its original form, this procedure can be iterated until convergence of $g_{i}^{*}$ and $h_{i}^{*}$.
 As shown in Ref.
-(leggett_spinboson), in the Ohmic regime ($s=1$), $h_{i}^{*}$ and $g_{i}^{*}$ converge to finite values $h_{i}(p\epsilon_{h}h_{i}/\omega_{c})^{\alpha/\left(1-\alpha\right)}$ and $g_{i}(pg_{i}/\omega_{c})^{2\alpha/\left(1-2\alpha\right)}$ if $\alpha_{i}<1$ and <1/2<1/2, respectively, otherwise, they converge to vanishing values, and tunneling ceases.
+[33], in the Ohmic regime ($s=1$), $h_{i}^{*}$ and $g_{i}^{*}$ converge to finite values $h_{i}(p\epsilon_{h}h_{i}/\omega_{c})^{\alpha/\left(1-\alpha\right)}$ and $g_{i}(pg_{i}/\omega_{c})^{2\alpha/\left(1-2\alpha\right)}$ if $\alpha_{i}<1$ and <1/2<1/2, respectively, otherwise, they converge to vanishing values, and tunneling ceases.
 For super-Ohmic dissipation ($s>1$), $g_{i}^{*}$ and $h_{i}^{*}$ are always finite.
@@
 However, near a phase transition, spins at different sites become correlated and, as shown in Ref.
-(smeared_hoyos_1), the effective dissipation parameter can change under RG.
+[5], the effective dissipation parameter can change under RG.
 The central question then becomes the following: under what conditions does the dissipation parameter flow to larger values?
-As we show in the following, this happens only when FM-like RRs appear, and in the FM phase.
-Surprisingly, not when PROD-like RRs are formed.
+As we show in the following, this happens only when FM-like RRs appear, and in the FM phase - surprisingly, not when PROD-like RRs are formed.
 V Strong-disorder renormalization group
@@
 We follow the methods of the dissipationless case in Ref.
-(2014_ashkinteller) and include dissipation using the strategy of Ref.
-(smeared_hoyos_2) which is the following.
+[30] and include dissipation using the strategy of Ref.
+[6] which is the following.
 The renormalization-group decimations are of two types: global and local, and are performed simultaneously when lowering the cutoff energy scale of the system.
@@
 IV).
-The local decimation concerns the coarse-grainning of a high-energy localized mode in a disordered system and, thus, is identical to the dissipationless case.
+The local decimation concerns the coarse grainning of a high-energy localized mode in a disordered system and, thus, is identical to the dissipationless case.
 Let $\Omega=\mathrm{max}\left\{K_{i},g_{i},J_{i},h_{i},\omega_{c}/p\right\}$ be the largest energy scale in the renormalized system. The decimation procedure now unfolds in the following steps:
+V.1 Global decimation
 (i) In lowering the energy scale from $\Omega$ to $\Omega-d\Omega$, we integrate out all the oscillators in the interval $\omega_{c}-pd\Omega<\omega_{k,i}<\omega_{c}=p\Omega$.
@@
 Notice that this is the global decimation procedure mentioned above.
-In contrast, the remaining steps are local decimations and are identical to the dissipationless case, except when it comes to renormalize the dissipation parameter $\alpha_{i}$.
-(ii) Let the two largest energy scales be $g_{i}$ and $K_{i}$.
-Thus, the unperturbed Hamiltonian is $H_{0}=-K_{i}\sigma_{i}^{z}\sigma_{i+1}^{z}-g{}_{i}\eta_{i}^{x}$ and all other terms in the Hamiltonian (4) containing $\boldsymbol{\sigma}_{i}$, $\boldsymbol{\sigma}_{i+1}$, and $\boldsymbol{\eta}_{i}$ are treated perturbatively up to second order.
+In contrast, the remaining steps are local decimations and are identical to the dissipationless case, except when it comes to renormalizing the dissipation parameter $\alpha_{i}$.
+V.2 Local decimations
+An interesting feature of the local decimations in the Ashkin-Teller model, when compared to the Ising model, is that there are two coupled dynamical variables in the problem: $\sigma$ and $\eta$ [see Eq.
+(4)].
+At zeroth order, they are not locally coupled and, therefore, the decimation must take into account two local energy scales: one for each variable.
+This is clear when one considers two decoupled Ising models.
+A local decimation must take into account the largest energy scale in each model.
+(ii) Let the largest energy scale in the system be $g_{i}$ and the second largest in that local cluster be $K_{i}$ (or vice versa).
+Thus the unperturbed Hamiltonian is $H_{0}=-K_{i}\sigma_{i}^{z}\sigma_{i+1}^{z}-g{}_{i}\eta_{i}^{x}$ and all other terms in the Hamiltonian (4) containing $\boldsymbol{\sigma}_{i}$, $\boldsymbol{\sigma}_{i+1}$, and $\boldsymbol{\eta}_{i}$ are treated perturbatively up to second order.
 As a result, the new effective Hamiltonian has the same form as the Hamiltonian (4) but with one fewer site and two fewer spin-1/2 degrees of freedom.
 The degrees of freedom $\boldsymbol{\sigma}_{i}$ and $\boldsymbol{\sigma}_{i+1}$ are clustered together and replaced by an effective spin-1/2 degree of freedom $\tilde{\boldsymbol{\sigma}}$ upon which the effective field is $\tilde{h}$.
-In addition, $\boldsymbol{\eta}_{i}$ becomes locked in the $x$-direction and does not contribute to the $z$-magnetization.
+In addition, $\boldsymbol{\eta}_{i}$ becomes locked in the $x$ direction and does not contribute to the $z$ magnetization.
 The effective coupling between the neighboring $\boldsymbol{\eta}$’s is $\tilde{J}$.
@@
 2(a), and the renormalized couplings are
-$\tilde{J}=\frac{2J_{i-1}J_{i}}{\tilde{g}_{i}},\quad\tilde{h}=\frac{2h_{i}h_{i+1}}{K_{i}}.$ (20)
-Now, we need to compute how the dissipation strength $\alpha$ changes under this decimation. Recalling that
+$\tilde{J}=\frac{2J_{i-1}J_{i}}{\tilde{g}_{i}}\mbox{ and }\tilde{h}=\frac{2h_{i}h_{i+1}}{K_{i}}.$ (21)
+Now we need to compute how the dissipation strength $\alpha$ changes under this decimation. We recall that
 $H_{\text{SB}}=$ $-\eta_{i}^{z}\sum_{k}\lambda_{k,i}x_{1;k,i}-\sigma_{i}^{z}\eta_{i}^{z}\sum_{k}\lambda_{k,i}x_{2;k,i}$
 $-\sigma_{i+1}^{z}\eta_{i+1}^{z}\sum_{k}\lambda_{k,i+1}x_{2;k,i+1},$ (22)
-and that the doubly degenerated ground state is $\left\{\left|\rightarrow_{i};\uparrow_{i}\uparrow_{i+1}\right\rangle,\left|\rightarrow_{i};\downarrow_{i}\downarrow_{i+1}\right\rangle\right\}$.
+and that the doubly degenerate ground state is $\left\{\left|\rightarrow_{i};\uparrow_{i}\uparrow_{i+1}\right\rangle,\left|\rightarrow_{i};\downarrow_{i}\downarrow_{i+1}\right\rangle\right\}$.
 Then, projecting $H_{\text{SB}}$ onto these states, we find that
@@
 IV).
-Thus, only the last term in (21) needs to be projected onto the ground-state manifold.
-Importantly, there is no change in the spectral function of the remaining baths, and, therefore, $\alpha_{i+1}$ remains unchanged.
-Figure 2: Decimation procedure when the two largest local energy scales in the system are (a) $g_{2}$ and $K_{2}$, (b) $g_{2}$ and $h_{2}$, and (c) $K_{2}$ and $J_{2}$ [in the main text, these correspond to the decimation cases (ii), (iii) and (iv), respectively].
-The couplings to the oscillator baths also renormalized and are not illustrated here.
-(iii) Let the two largest energy scales be $g_{i}$ and $h_{i}$.
+Thus only the last term in (22) needs to be projected onto the ground-state manifold.
+Importantly, there is no change in the spectral function of the remaining baths and, therefore, $\alpha_{i+1}$ remains unchanged.
+Figure 2: Decimation procedure: (a) Global decimation due to the coupling with the dissipative bath [with $\tilde{h}_{i}$ and $\tilde{g}_{i}$ given by (19) and (20), respectively].
+(b) Local decimations due to a broad distribution of local energy scales.
+When the two largest local energy scales in the system are (i) $g_{2}$ and $K_{2}$ [with $\tilde{J}$ and $\tilde{h}$ given by (21)], (ii) $g_{2}$ and $h_{2}$ [with $\tilde{K}$ and $\tilde{J}$ given by (24)], and (iii) $K_{2}$ and $J_{2}$ [with $\tilde{g}$ and $\tilde{h}$ given by (25)] In the main text, these correspond to the decimation cases (i), (ii), (iii) and (iv), respectively.
+The couplings between the spin clusters and the oscillator baths are also renormalized and are not illustrated here.
+(iii) Let the largest energy scales be $g_{i}$ and the second largest in that local cluster be $h_{i}$.
 In this case, all tunneling frequencies have fully (or nearly) converged to their fixed-point values $g^{*}$ and $h^{*}$ in the adiabatic renormalization.
@@
 2(b) and the renormalized quantities are
-$\tilde{K}=\frac{K_{i-1}K_{i}}{2h_{i}},\quad\tilde{J}=\frac{J_{i-1}J_{i}}{g_{i}+h_{i}}.$ (23)
+$\tilde{K}=\frac{K_{i-1}K_{i}}{2h_{i}}\mbox{ and }\tilde{J}=\frac{J_{i-1}J_{i}}{g_{i}+h_{i}}.$ (24)
 Since no spin clusters are formed in this step, the projection of $H_{\text{SB}}$ into the ground state results in zero.
 Hence, $\alpha$ remains unchanged (except, evidently, for those ones decimated).
-(iv) The final case to consider is when $K_{i}$ and $J_{i}$ are the largest energy scales.
+(iv) The final case to consider is when $K_{i}$ is the largest energy scale in the system and $J_{i}$ is the second largest in that local cluster.
 The unperturbed Hamiltonian is $H_{0}=-K_{i}\sigma_{i}^{z}\sigma_{i+1}^{z}-J_{i}\eta_{i}^{z}\eta_{i+1}^{z}-J_{i}\sigma_{i}^{z}\sigma_{i+1}^{z}\eta_{i}^{z}\eta_{i+1}^{z}$.
@@
 2(c) where the renormalized quantities are
-$\tilde{g}=\frac{g_{i}g_{i+1}}{2J_{i}},\quad\tilde{h}=\frac{h_{i}h_{i+1}}{K_{i}+J_{i}}.$ (24)
-For this step, we need to take a closer look on what happens to the oscillator baths. The system-bath Hamiltonian is
+$\tilde{g}=\frac{g_{i}g_{i+1}}{2J_{i}}\mbox{ and }\tilde{h}=\frac{h_{i}h_{i+1}}{K_{i}+J_{i}}.$ (25)
+For this step, we need to take a closer look at what happens to the oscillator baths. The system-bath Hamiltonian is
 $H_{\text{SB}}=$ $-\eta_{i}^{z}\sum_{k}\lambda_{k,i}x_{1;k,i}-\sigma_{i}^{z}\eta_{i}^{z}\sum_{k}\lambda_{k,i}x_{2;k,i}$
 $-\eta_{i+1}^{z}\sum_{k}\lambda_{k,i+1}x_{1;k,i+1}-\sigma_{i+1}^{z}\eta_{i+1}^{z}\sum_{k}\lambda_{k,i+1}x_{2;k,i+1},$ (26)
-and needs to be projected onto the 4-fold-degenerate ground state of $H_{0}$, which is $\left\{\left|\eta;\sigma\right\rangle\right\}$ (where $\eta_{j}^{z}\left|\eta;\sigma\right\rangle=\eta\left|\eta;\sigma\right\rangle$, $\sigma_{j}^{z}\left|\eta;\sigma\right\rangle=\sigma\left|\eta;\sigma\right\rangle$, with $j=i,i+1$ and $\eta,\sigma=\pm 1$ representing the effective spins).
+and needs to be projected onto the fourfold-degenerate ground state of $H_{0}$, which is $\left\{\left|\eta;\sigma\right\rangle\right\}$ (where $\eta_{j}^{z}\left|\eta;\sigma\right\rangle=\eta\left|\eta;\sigma\right\rangle$, $\sigma_{j}^{z}\left|\eta;\sigma\right\rangle=\sigma\left|\eta;\sigma\right\rangle$, with $j=i,i+1$ and $\eta,\sigma=\pm 1$ representing the effective spins).
 Since $H_{\text{SB}}$ is diagonal in all four states, the projection can be easily performed, yielding the following effective Hamiltonian
 $\tilde{H}_{\text{SB}}=$ $-\tilde{\eta}^{z}\sum_{k,j}\lambda_{k,j}x_{1;k,j}-\tilde{\sigma}^{z}\tilde{\eta}^{z}\sum_{k,j}\lambda_{k,j}x_{2;k,j},$ (27)
-where the sum in the index $j$ run from $i$ to $i+1$. Consequently, the spectral density of the local renormalized bath is
+where the sum in the index $j$ runs from $i$ to $i+1$. Consequently, the spectral density of the local renormalized bath is
 $\tilde{F}(\omega)=\frac{\pi}{2}\sum_{k,j}\frac{\lambda_{k,j}^{2}}{m_{k,j}\omega_{k,j}}\delta(\omega-\omega_{k,j})=\frac{\pi}{2}\tilde{\alpha}\frac{\omega^{s}}{\omega_{I}^{s-1}}e^{-\omega/\omega_{c}},$ (28)
@@
 (29), it becomes clear that the effective dissipation strength increases as the volume of the FM cluster, just like in the case of a direct transition between the PM and FM phases studied in Refs.
-(smeared_hoyos_1; smeared_hoyos_2).
-In addition, it is possible to relate the dissipation strength with the effective magnetic moment as both increases with the cluster volume.
-It is a convenient relation when analyzing the SDRG flow equations for the distributions of couplings, fields and dissipation parameters (smeared_hoyos_2).
-VI Phase diagram
+[5, 6].
+In addition, it is possible to relate the dissipation strength with the effective magnetic moment as both increase with the cluster volume.
+It is a convenient relation when analyzing the SDRG flow equations for the distributions of couplings, fields and dissipation parameters [6].
+VI Phase diagram: dissipative case
 We are now able to describe the phase diagram of our model Hamiltonian (5) sketched in Fig.
@@
 However, they do cover all decimations necessary to analyze the PM-PROD and PROD-FM transitions as shown in Ref.
-(2014_ashkinteller).
-To review the undamped case, we start noticing that the PM-PROD transition happens from the competition between the transverse field $h$ and the 4-spin coupling $K$ [see the first sum in the Hamiltonian (4), the other terms are just bystanders].
+[30].
+The missing decimations are those where $J_{i}$ or $h_{i}$ is the largest energy scale.
+They are important in analyzing the $\epsilon<1$ case where the two chains decouple and $\epsilon$ renormalizes to zero [29].
+This situation, thus, was analyzed elsewhere with and without dissipation [29, 5, 6]
+VI.1 Reviewing the dissipationless case
+To briefly review the undamped case (see further details in Sec.
+II.3), we start noticing that the PM-PROD transition happens from the competition between the transverse field $h$ and the 4-spin coupling $K$ [see the first sum in the Hamiltonian (4); the other terms are just bystanders].
 The relevant decimations are, thus, those described in cases (ii) and (iii) in the previous section.
@@
 (21) and (24), the transition line happens when the typical values are $2h_{\text{typ}}=K_{\text{typ}}$.
-Thus, the transition line in the undamped case is at $\delta\approx\epsilon_{I}/2$.
+Thus the transition line in the undamped case is at $\delta\approx\epsilon_{I}/2$.
 For the PROD-FM transition, the competition is between the 2-spin couplings $J$ and $g$ [see the second sum in the Hamiltonian (4)].
@@
 Following the same reasoning as before, the transition line is at $\delta\approx 2/\epsilon_{I}.$
+VI.2 Dissipative case: Sharp transition
 Let us now analyze the effects of weak Ohmic dissipation $\alpha\ll 1$ and $s=1$ in Eq.
@@
 (19)].
-Therefore, Ohmic dissipation is a irrelevant perturbation to the PM-PROD quantum phase transition.
+Therefore, Ohmic dissipation is an irrelevant perturbation to the PM-PROD quantum phase transition.
 Its only effect is a slight change of the critical line.
-This same conclusion applies to the case of super-Ohmic dissipation ($s>1$) which is qualitatively weaker when compared to Ohmic dissipation (leggett_spinboson).
+This same conclusion applies to the case of super-Ohmic dissipation ($s>1$) which is qualitatively weaker when compared to Ohmic dissipation [33].
 Very likely, the same conclusion may also apply to the case of sub-Ohmic dissipation ($s<1$).
-The effects of dissipation is much more dramatic in the PROD-FM transition.
+VI.3 Dissipative case: Smeared transition
+The effects of dissipation are much more dramatic in the PROD-FM transition.
 As pointed out in Secs.
 II and V, FM-like RRs are formed around that transition.
-This is because a local FM coupling $J$ can be one of the strongest energy scale in the system [decimation case (iv) in Sec.
+This is because a local FM coupling $J$ can be the second largest energy scale in the local cluster [decimation case (iv) in Sec.
 V].
@@
 (28) and (29)].
-Thus, a sufficiently large RR (and they can be arbitrarily large in a Griffiths phase (fisher_2)) have an effective dissipation parameter $\tilde{\alpha}>\alpha_{c}=1/2$.
+Thus, a sufficiently large RR (and they can be arbitrarily large in a Griffiths phase [2]) has an effective dissipation parameter $\tilde{\alpha}>\alpha_{c}=1/2$.
 As a consequence, these larger RRs will cease tunneling ($\tilde{g}\rightarrow 0$) and undergo their own FM transition when Ohmic dissipation is considered.
@@
 Any residual interaction between these larger RRs will align them together and the total magnetization will not average out.
-The Griffiths PROD phase of the undamped system now becomes a inhomogeneous FM phase due to Ohmic dissipation.
-Upon lowering $\delta$, this inhomogeneous FM phase crosses over to a weakly ordered phase (the former Griffiths FM phase of the undamped system) and, finally, crosses over to the conventional FM phase when further lowering $\delta.$ It is important to notice that this phenomenon is akin to the smearing in the PM-FM transition in the random transverse-field Ising chain studied in Refs.
-(smeared_hoyos_1; smeared_hoyos_2).
+The Griffiths PROD phase of the undamped system now becomes an inhomogeneous FM phase due to Ohmic dissipation.
+Upon lowering $\delta$, this inhomogeneous FM phase crosses over to a weakly ordered phase (the former Griffiths FM phase of the undamped system) and, finally, crosses over to the conventional FM phase when further lowering $\delta.$ Having the different names for the IFM and WO phases is to emphasize that in the former, the magnetization appears only in some frozen and large rare regions while in the latter, the magnetization happens in the bulk.
+It is important to notice that this phenomenon is akin to the smearing in the PM-FM transition in the random transverse-field Ising chain studied in Refs.
+[5, 6].
 Thus, both PM-FM ($\epsilon_{I}<1$) and PROD-FM ($\epsilon_{I}>1$) transitions become smeared in the presence of Ohmic dissipation as sketched in Fig.
@@
 For weak super-Ohmic dissipation, on the other hand, all transitions remain sharp because a RR cannot stop tunneling due to the coupling to the oscillator bath regardless the effective value of $\tilde{\alpha}$.
+VI.4 Observables
 It is interesting to quantify some key observables in those transitions.
@@
 We refer the reader to Ref.
-(2014_ashkinteller) for the critical behavior of many observables.
+[30] for the critical behavior of many observables.
 For the smearing transitions, many observables were quantified in Refs.
-(smeared_hoyos_1; smeared_hoyos_2).
+[5, 6].
 Here, we only quote the magnetization at the onset of the inhomogeneous FM phase: $\overline{\left\langle\sigma_{i}^{z}\right\rangle}=\overline{\left\langle S_{1,i}^{z}\right\rangle}=\overline{\left\langle S_{2,i}^{z}\right\rangle}\sim w^{1/\alpha},$ where $\overline{\left\langle\cdot\right\rangle}$ denotes the disorder and quantum averages.
@@
 Notice that $w$ plays the role of distance from transition.
-VII Discussion and conclusions
+VII Discussions and conclusions
 In this work, we investigated the effects of dissipation in the phase transitions of the quantum Ashkin-Teller chain in the presence of quenched disorder.
-The dissipationless model exhibits three quantum phase transitions (see Fig.
-1(a)): a FM-PM transition, a FM-PROD transition, and a PM-PROD transtion.
-All belonging to the universality class of the random transverse-field Ising chain and displaying off-critical quantum Griffiths phases.
+The dissipationless model exhibits three quantum phase transitions [see Fig.
+1(a)]: a FM-PM transition, a FM-PROD transition, and a PM-PROD transition, all belonging to the universality class of the random transverse-field Ising chain and displaying off-critical quantum Griffiths phases.
 Introducing dissipation of the quantum fluctuations, surprisingly, does not smear all of these phase transitions.
-Only the transitions involving the FM phase are smeared (see Fig.
-1(b)).
-In the more familiar model of the random-transverse field Ising chain, the low-energy physics is dominated by the so-called rare regions in the Griffiths phases.
+Only the transitions involving the FM phase are smeared [see Fig.
+1(b)].
+In the more familiar model of the random transverse-field Ising chain, the low-energy physics is dominated by the so-called rare regions in the Griffiths phases.
 These are regions which are locally FM and weakly coupled to the bulk.
@@
 Smearing due to dissipation occurs because that slow dynamics is completely halted by Ohmic and sub-Ohmic dissipation.
-The coherent tunneling requires a coherent changing in all oscillators rest position.
-When that set of oscillators is sufficiently slow, the rare region stop tunneling and develops its own magnetic order regardless the bulk.
+The coherent tunneling requires a coherent changing in all oscillators’ rest position.
+When that set of oscillators is sufficiently slow, the rare region stops tunneling and develops its own magnetic order regardless of the bulk.
 It turns out that the slow tunneling of the rare regions in the PM-PROD phase does not require coherent tunneling of all their constituents.
 This is because the magnetic order of the PROD phase is a composite one: each color spin is disordered, $\left\langle S_{1}^{z}\right\rangle=\left\langle S_{2}^{z}\right\rangle=0$, while the composite magnetic order finite, $\left\langle S_{1}^{z}S_{2}^{z}\right\rangle\neq 0$.
-Thus, dissipation do not couple directly to the PROD order parameter and, therefore, the transition remains sharp.
+Thus dissipation does not couple directly to the PROD order parameter and, therefore, the transition remains sharp.
 It is interesting to consider whether similar situations appear in other contexts.
-Composite (quadrupole-like) nematic order appears in diverse contexts such as in frustrated magnets (as, for instance, in the context of order-by-disorder in the $J_{1}$-$J_{2}$ Heisenberg model or in easy-plane pyrochlores) and in iron pnictides.
-Giving the results here uncovered, dissipation due to the coupling to gapless modes (such as that of a metallic Fermi liquid) may not couple to that composite magnetic order, and, consequently, the fluctuations of the associated nematic rare regions may not be overdamped by dissipation.
+Composite (quadrupolelike) nematic order appears in diverse contexts such as in frustrated magnets (as, for instance, in the context of order by disorder in the $J_{1}$-$J_{2}$ Heisenberg model or in easy plane pyrochlores) and in iron pnictides.
+Giving the results here uncovered that dissipation due to the coupling to gapless modes (such as that of a metallic Fermi liquid) may not couple to that composite magnetic order, and, consequently, the fluctuations of the associated nematic rare regions may not be overdamped by dissipation.
 It is also interesting to test our result in other models.
-In Appendix A we consider two possible modifications on the dissipation of the quantum fluctuations: one in which the couplings to the oscillator bath is color-dependent and another in which there is only one oscillator bath for both color spins.
+In the Appendix A we consider two possible modifications on the dissipation of the quantum fluctuations: one in which the couplings to the oscillator bath are color dependent and another in which there is only one oscillator bath for both color spins.
 In both cases, our result remains: the smearing occurs only to the transitions to the FM phase.
 Finally, we argue that our results are not restricted to one spatial dimension.
-Our SDRG method can be applied to any dimension and the smearing (or lack of) will not depend on the underling coordination number.
+Our SDRG method can be applied to any dimension and the smearing (or lack thereof) will not depend on the underlying coordination number.
 Acknowledgements.
@@
 To check how dissipation will affect the phase transitions, we project the coupling Hamiltonian $H_{\text{SB}}$ onto the ground states of each decimation case.
-For cases (ii) and (iii), the coupling Hamiltonian have the same form as the color-independent case (apart from the color dependence of the $\lambda$’s).
+For cases (ii) and (iii), the coupling Hamiltonian has the same form as the color-independent case (apart from the color dependence of the $\lambda$’s).
 Therefore, there is no renormalization of the dissipation strength $\alpha$.
@@
 A.2 Single dissipation bath
-Here, we investigate whether our results on the model Hamiltonian (5) also applies to another related dissipative model.
-In (5), each spin is coupled to its own set of oscillator bath.
+Here, we investigate whether our results on the model Hamiltonian (5) also apply to another related dissipative model.
+In (5), each spin is coupled to its own set of oscillator baths.
 We now consider the case in which different color-spins are coupled to the same set of oscillators.
@@
 $\triangle V=\frac{1}{2}\sum_{k,i}m_{k,i}\omega_{k,i}^{2}x_{0;k,i}^{2}\left(1+\sigma_{i}^{z}\right)^{2},$
-where $x_{0;k,i}=\lambda_{k,i}/m_{k,i}\omega_{k,i}^{2}$ is the counter-term (caldeiraleggett).
+where $x_{0;k,i}=\lambda_{k,i}/m_{k,i}\omega_{k,i}^{2}$ is the counterterm [32].
 It must be added if we wish that the coupling to the bath introduces only dissipation, and not effective fields [which appears when completing the square in the last line of Eq.
@@
 $=$ $\sum_{\omega_{k,i}>\omega_{l}}\left(\frac{p_{k,i}^{2}}{2m_{k,i}}+\frac{m_{k,i}\omega_{k,i}^{2}\left(x_{k,i}+x_{0k,i}\eta_{i}^{z}(1+\sigma_{i}^{z})\right)^{2}}{2}\right),$ (36)
-and treat the tunneling terms of the $i$-th site in first order of perturbation theory.
+and treat the tunneling terms of the $i$th site in first order of perturbation theory.
 The definition of $\omega_{l}$ is the same as in Sec.
 IV.
-The 4-fold degenerated ground state of $H_{0}$ is given by
-$\left|\Psi_{++}\right\rangle=$ $\left|+,+\right\rangle\prod_{\omega_{k,i}>\omega_{l}}\left|G_{k,i;++}\right\rangle,$
-$\left|\Psi_{-+}\right\rangle=$ $\left|-,+\right\rangle\prod_{\omega_{k,i}>\omega_{l}}\left|G_{k,i;-+}\right\rangle,$ (36)
-$\left|\Psi_{\pm-}\right\rangle=$ $\left|\pm,-\right\rangle\prod_{\omega_{k,i}>\omega_{l}}\left|0_{k,i}\right\rangle,$
-where we are using the same notation as Sec. IV (first and indices to $\eta$ and $\sigma$ variables, respectively). The oscillators’ ground states are
+The fourfold degenerate ground state of $H_{0}$ is given by
+$\left|\Psi_{++}\right\rangle=$ $\left|+,+\right\rangle\bigotimes_{\omega_{k,i}>\omega_{l}}\left|G_{k,i;++}\right\rangle,$
+$\left|\Psi_{-+}\right\rangle=$ $\left|-,+\right\rangle\bigotimes_{\omega_{k,i}>\omega_{l}}\left|G_{k,i;-+}\right\rangle,$ (37)
+$\left|\Psi_{\pm-}\right\rangle=$ $\left|\pm,-\right\rangle\bigotimes_{\omega_{k,i}>\omega_{l}}\left|0_{k,i}\right\rangle,$
+where we are using the same notation as Sec. IV (first and second indices to $\eta$ and $\sigma$ variables, respectively). The oscillators’ ground states are
 $\left|G_{k,i;\eta_{i}^{z}\sigma_{i}^{z}}\right\rangle=e^{2i\eta_{i}^{z}\sigma_{i}^{z}p_{k,i}x_{0;k,i}}\left|0_{k,i}\right\rangle,$ (38)
 and $\left|0_{k,i}\right\rangle$ is the ground state of the oscillator with the rest position at the origin. Projecting the perturbation
-$H_{1}=-h_{i}\sigma_{i}^{x}(1+\eta_{i}^{x})-g_{i}\eta_{i}^{x},$ (38)
+$H_{1}=-h_{i}\sigma_{i}^{x}(1+\eta_{i}^{x})-g_{i}\eta_{i}^{x}$ (39)
 onto the ground state manyfold (the corresponding matrix element being $H_{\text{eff}\eta^{z}\sigma^{z},\eta^{\prime z}\sigma^{\prime z}}=\left\langle\Psi_{\eta^{z}\sigma^{z}}\right|H_{1}\left|\Psi_{\eta^{\prime z}\sigma^{\prime z}}\right\rangle$), we find the effective Hamiltonian
 $H_{\text{eff}}=$ $-\tilde{h_{i}}\sigma_{i}^{x}\left(1+\eta_{i}^{x}\right)-g_{i}\eta_{i}^{x}\left(\frac{1-\sigma_{i}^{z}}{2}\right)$
-$-\tilde{g_{i}}\eta_{i}^{x}\left(\frac{1+\sigma_{i}^{z}}{2}\right)$ (39)
+$-\tilde{g_{i}}\eta_{i}^{x}\left(\frac{1+\sigma_{i}^{z}}{2}\right),$ (40)
 with the renormalized tunneling frequencies
@@
 From Eq.
-(39), it follows that the tunneling frequency $g_{i}$ splits in two parts: one that decreases under adiabatic renormalization and other that remains unaffected.
+(40), it follows that the tunneling frequency $g_{i}$ splits into two parts: one that decreases under adiabatic renormalization and another that remains unaffected.
 This occurs because both states $\left|\Psi_{\pm-}\right\rangle$ are centered at the same position, such that the tunneling between states $\left|\Psi_{+-}\right\rangle$ and $\left|\Psi_{--}\right\rangle$, driven by the term $g_{i}\eta_{i}^{x}$ in $H_{1}$, does not change the equilibrium position of the oscillators.